### v1

# Labels Matter More Than Models: Quantifying the Benefit of Supervised Time Series Anomaly Detection

Abstract
Time series anomaly detection (TSAD) is a critical data mining task often constrained by label scarcity. Consequently, current research predominantly focuses on Unsupervised Time-series Anomaly Detection (UTAD), relying on complex architectures to model normal data distributions. However, this approach often overlooks the significant performance gains available from limited anomaly labels achievable in practical scenarios. This paper challenges the premise that architectural complexity is the optimal path for TSAD. We conduct the first methodical comparison between supervised and unsupervised paradigms and introduce STAND, a streamlined supervised baseline. Extensive experiments on five public datasets demonstrate that: (1) Labels matter more than models: under a limited labeling budget, simple supervised models significantly outperform complex state-of-the-art unsupervised methods; (2) Supervision yields higher returns: the performance gain from minimal supervision far exceeds that from architectural innovations; and (3) Practicality: STAND exhibits superior prediction consistency and anomaly localization compared to unsupervised counterparts. These findings advocate for a data-centric shift in TSAD research, emphasizing label utilization over purely algorithmic complexity. The code is publicly available at https://github.com/EmorZz1G/STAND.
I Introduction
Time series anomaly detection (TSAD) is a crucial and challenging task in time series data mining, with broad applications in fields such as industrial system monitoring, cybersecurity, and health surveillance [adamembls, cce, simad, TSB-UAD].
Due to the scarcity of anomaly samples and the high cost of labeling in TSAD, unsupervised time series anomaly detection (UTAD) methods have garnered significant attention in recent years [bls, simad, catch, M2N2]. Typically, unsupervised methods assume that the training time series data primarily consists of normal samples. They detect anomalies by learning the patterns of these normal samples, and this approach currently represents the mainstream of research. Because these methods introduce the prior assumption that the training data is composed entirely of normal samples, they are also referred to as semi-supervised techniques in some literature [TSB-UAD, catch, ppt]. Moving forward, we will denote this specific category as Unsupervised Time series Anomaly Detection Type II (UTAD-II), as illustrated in Figure 1(b).
Conversely, earlier studies trained and detected anomalies using unlabeled datasets that might contain anomaly samples, leading to their original designation as unsupervised techniques [TSB-UAD, pca]. We will refer to this category as Unsupervised Time series Anomaly Detection Type I (UTAD-I), as shown in Figure 1(a).
Although there may be some debate regarding the precise categorization of unsupervised TSAD, the fundamental consensus in current research is that the training phase does not directly use anomaly sample labels for model optimization. Unless otherwise specified, this paper adopts this overarching philosophy and collectively refers to both UTAD-I and UTAD-II as unsupervised techniques (UTAD).
In contrast, Supervised Time series Anomaly Detection (STAD) methods utilize anomaly sample labels during the training phase to optimize performance (as shown in Figure 1(c)), but such methods have received limited research attention [STSAD, ppt, label2].
Figure 1: Classification of time series anomaly detection approaches.
However, in practical applications, a small amount of anomaly sample labels can be acquired, and the purely unsupervised assumption often struggles to hold true [label1]. Existing studies largely overlook the significance of utilizing these few labels in time series anomaly detection, focusing instead primarily on improvements to model architecture.
Is relentlessly pursuing complex model architecture optimization truly the best path for time series anomaly detection research? Given a limited labeling budget, how significant is the performance gain achieved by incorporating labels?
To address these critical questions, we propose a simple Supervised Time-series ANomaly Detection (STAND) baseline method. Furthermore, we present the first dedicated comparison between supervised and unsupervised methods to date. Through extensive experimentation, we derive the following key findings and contributions:
  1. 1.
    Labels Matter More Than Models: Under a limited labeling budget, the simple STAD method, STAND, significantly outperforms existing complex unsupervised time series anomaly detection methods. As illustrated in Figure 2, the STAD category consistently yields superior performance compared to both UTAD-I and UTAD-II.
  2. 2.
    The Benefit of Supervision: Given a limited labeling budget, the performance improvement gained from supervision far outweighs the incremental gains from model architecture improvements.
  3. 3.
    Readability & Practicality: Existing Unsupervised Time series Anomaly Detection (UTAD) methods often fail to clearly delineate or anchor anomaly segments; in contrast, STAD methods successfully detect the majority of anomalies while maintaining consistent predictions, demonstrating stronger practical applicability.
  4. 4.
    The First Supervised/Unsupervised TSAD Benchmark and Library: We present the first work to incorporate supervised methods into TSAD evaluation landscape. We established a complete evaluation pipeline and open-source our code [^1] STAND: https://github.com/EmorZz1G/STAND.
Figure 2: Multi-metric performance comparison across different categories of time series anomaly detection methods.
II Related Work
II-A Algorithm-Centric Time Series Anomaly Detection
Early unsupervised Time Series Anomaly Detection (UTAD) methods were primarily based on classical machine learning techniques, such as Principal Component Analysis (PCA) [pca], Local Outlier Factor (LOF) [ml_lof2000], and Isolation Forest (IForest) [ml_IsolationForest2008]. These methods transform time series data into static feature vectors using techniques like sliding window segmentation and statistical feature extraction, which are then fed into traditional anomaly detection algorithms. However, these approaches heavily depend on the quality of hand-crafted features, and their model structures are often insufficiently adapted to the dynamic nature of time series data.
With the advancement of deep learning technology, its powerful capabilities for automatic feature extraction and complex pattern modeling have been widely applied in the field of UTAD. Existing algorithm-centric research in this area can be broadly classified into two main technical paths:
Path I: Focusing on Model Architecture Improvement to Enhance Time Series Feature Extraction Capabilities. This includes methods based on improvements to classical neural networks [tcn_ed, usad, M2N2]; Transformer-based approaches [anom_trans], which leverage attention mechanisms to improve long-range dependency modeling; Mamba-based methods [s6_MemMambaAD, s6_wangMambaEffectiveTime2024, s6_behrouzChimeraEffectivelyModeling2024], which enhance the efficiency of processing long time sequences through state space models; and Large Language Model (LLM)-based methods [gpt2adpt, llm_AnomalyLLM2024], which utilize the general knowledge learned by LLMs for anomaly recognition.
Path II: Focusing on Establishing Anomaly Priors to Optimize Anomaly Determination. Examples include networks based on one-class classification and Random Forest principles [deep_svdd, deep_if], which introduce the concept of data centricity to identify anomalies; methods based on dissimilarity learning [patchad, simads, simad, dcdetector], which employ proxy strategies to learn the similarity distribution between normal and anomalous instances; and methods based on time-frequency information learning [TFAD, dual-tf, NPSR], which transform time series data into the frequency domain to capture multi-granularity anomalies.
All the aforementioned paths originate from the algorithmic level, attempting to overcome the label scarcity problem by constructing powerful models, but they fail to address the core contradiction directly from the data perspective.
II-B Data-Centric Time Series Anomaly Detection
Unlike the goal of algorithm-centric research, data-centric time series anomaly detection research revolves around optimizing the data layer to compensate for the deficiency of labels in unsupervised settings. However, due to the low prevalence of anomaly events and the high cost of labeling in most practical scenarios, studies that directly leverage high-quality anomaly labels to optimize models remain scarce [STSAD, sv_landauerTimeSeriesAnalysis2018, sv_belayUnsupervisedAnomalyDetection2023]. Based on this constraint, past data-centric research on time series anomaly detection has mainly explored two optimization paths:
Path I: Pseudo-Sample-Based Data Augmentation Strategies to Expand Data Value with Limited Information. The core of this path is to substitute real labels with data transformations, creating pseudo-anomaly samples or task-oriented samples to provide learning signals for the model. This is implemented in two ways: First, unsupervised pseudo-sample generation, where normal samples are modified into anomaly samples in a label-free setting by simulating anomaly mechanisms (e.g., anomaly injection, segment shuffling, etc.). This guides the model to explore anomaly patterns and enhance detection capability [couta, CutAddPaste, trs_jeongAnomalyBERTSelfSupervisedTransformer2023]. Second, construction of task-guided samples, where, for instance, noisy time series data are treated as boundary samples between normal and anomalous states, or similar samples are generated via operations like Mixup, interpolation, or inversion to assist with model initialization [adamembls, amsl, simad, s6_zengCMambaChannelCorrelation2024].
Path II: New Sample Generation Strategies Based on Distribution Modeling, Guiding Pattern Learning through Data Generation. This path utilizes generative models (e.g., GANs, VAEs, Diffusion Models) to learn the distribution characteristics of normal time series data, thereby generating new samples that conform to realistic patterns and reinforcing the model’s perception of normal modes [GenIAS, app_liDiffTADDenoisingDiffusion2024a, gan_duGANBasedAnomalyDetection2021]. The generated samples can directly augment the training set, mitigating data scarcity. Simultaneously, an anomaly determination baseline is established through the detection of deviations from the learned normal distribution, forming a closed loop between data generation and anomaly detection.
Despite the fact that these studies have explored the improvement of model performance via data quality enhancement, their core focus remains restricted to the unsupervised framework. They essentially bypass the reliance on true labels through pseudo-samples or generated samples, failing to fully exploit the few anomaly labels that are practically available. This represents a shared limitation of both the algorithm-centric and current data-centric approaches: both inherently assume a completely label-free prerequisite, yet overlook the potential gain in performance from the supervisory signal provided by limited labeling.
III Methodology
To clearly contrast the core differences between various paradigms of TSAD and to highlight the value of the supervisory signal, this section first establishes the mathematical definitions for unsupervised, supervised, and our proposed STAND methods. Following this, the specific implementation of the STAND baseline is detailed.
III-A Definitions
Definition 1 (Time Series Anomaly Detection).
The multivariate time series sample is represented as $\mathbf{X}=[x_{1},x_{2},\dots,x_{T}]^{\top}\in\mathbb{R}^{T\times C}$, where $T$ is the time series length, $C$ denotes the number of variables, and $x_{t}\in\mathbb{R}^{C}$ is the multivariate observation at time $t$. The corresponding time-step labels are $\mathbf{y}=[y_{1},y_{2},\dots,y_{T}]^{\top}$, where $y_{t}=0$ indicates a normal time step and $y_{t}=1$ indicates an anomalous time step. The objective of anomaly detection is to learn a mapping function $f:\mathbb{R}^{T}\rightarrow\mathbb{R}$, which outputs the anomaly score for each time step, $\mathbf{S}=f(\mathbf{X})$. A preset threshold $\tau$ is then used for class determination: if $s_{t}>\tau$, time step $t$ is classified as anomalous; otherwise, it is normal.
Definition 2 (Unsupervised Time-series Anomaly Detection (UTAD)).
The core assumption of UTAD is that the proportion of anomaly samples in the training data is extremely low or the training data is entirely anomaly-free. The training set is defined as $\mathcal{D}_{\text{trn}}=\{x_{i}\}_{i=1}^{N_{\text{trn}}}$. The method’s goal is to capture the distribution characteristics of normal time series via unsupervised learning, generating training set anomaly scores $\mathbf{S}_{\text{trn}}=f(\mathbf{X}_{\text{trn}})$, and then inferring the scores $\mathbf{S}_{\text{test}}=f(\mathbf{X}_{\text{test}})$ on the test set $\mathcal{D}_{\text{test}}=\{x_{j}\}_{j=1}^{N_{\text{test}}}$, which contains anomalies.
Definition 3 (Supervised Time-series Anomaly Detection (STAD)).
STAD is applicable in scenarios where a small number of anomaly labels are available. Its training set is defined as the labeled data $\mathcal{D}_{\text{trn}}=\{(x_{i},y_{i})\}_{i=1}^{N_{\text{trn}}}$. The method’s goal is to leverage supervisory information to guide the model in learning the mapping relationship, generating training set anomaly scores $\mathbf{S}_{\text{trn}}=f(\mathbf{X}_{\text{trn}},\mathbf{y})$, and subsequently detecting anomalies in the test set $\mathcal{D}_{\text{test}}$.
III-B Proposed Method
In this section, we introduce our proposed Supervised Time-series ANomaly Detection (STAND) framework. Our core hypothesis is that even a simple model, when guided by a small amount of supervision, can outperform complex unsupervised architectures. Consequently, STAND is intentionally designed with a simple yet effective architecture to serve as a strong baseline and empirically validate the significant impact of labels. The framework treats time-series anomaly detection as a sequence-to-sequence binary classification task at the timestamp level. As illustrated in Figure 3, STAND is comprised of three main components: a Feature Embedding, a Temporal Encoding Module, and an Anomaly Scoring Module.
Figure 3: The overall architecture of the STAND  framework. An input time series is first projected into latent space by the feature embedding. The resulting latent sequence is then fed into the temporal encoding module, which captures bidirectional temporal context. Finally, an anomaly scoring module outputs anomaly scores for each time step.
III-B1 Model Architecture
Feature Embedding. The raw input at each timestamp, $x_{t}\in\mathbb{R}^{C}$, often resides in a feature space that may not be optimal for separating normal and anomalous patterns. To address this, we first employ a Feature Embedding to project the input features into a more discriminative latent space. This module enhances the model’s capacity to learn complex, non-linear relationships between the input variables before they are processed sequentially.
As defined in our implementation, this module is a simple Multi-Layer Perceptron (MLP) with two linear layers, each followed by a Gaussian Error Linear Unit (GELU) activation function [gelu] and Layer Normalization [layernorm]. The process for a single timestamp $x_{t}$ can be formulated as:
$h_{t}^{(0)}=x_{t}$ (1)
$h_{t}^{(l)}=\text{LayerNorm}(\text{GELU}(\mathbf{W}^{(l)}h_{t}^{(l-1)}+b^{(l)}))$ (2)
where $l$ is the layer index in the MLP. The final output of this module is a sequence of embeddings $\mathbf{H}^{e}=[h_{1}^{e},h_{2}^{e},\dots,h_{T}^{e}]^{\top}\in\mathbb{R}^{T\times d_{\text{model}}}$, where $d_{\text{model}}$ is the hidden dimension of the model. This embedding process is applied independently to each timestamp. An option to bypass this module is included for ablation studies, in which case the input is fed directly to the next module.
Temporal Encoding Module. To capture the temporal dependencies inherent in time-series data, the sequence of embeddings $\mathbf{H}^{e}$ is passed to a Temporal Encoding Module. The choice of encoder is critical for modeling the contextual information that distinguishes an anomaly from normal behavior. We utilize a Long Short-Term Memory (LSTM) network [lstm], a proven and effective recurrent neural network (RNN) architecture for sequential data. LSTMs use a gating mechanism (input, forget, and output gates) to regulate information flow, enabling them to capture long-range dependencies while mitigating the vanishing gradient problem.
To provide the model with a comprehensive contextual understanding for each timestamp, we employ a bidirectional LSTM. A bidirectional LSTM processes the input sequence in both forward (from $t=1$ to $T$) and backward (from $t=T$ to $1$) directions. The hidden state for each timestamp $t$ is the concatenation of the forward hidden state $\overrightarrow{h_{t}}$ and the backward hidden state $\overleftarrow{h_{t}}$:
$\overrightarrow{h_{t}}=\text{LSTM}_{\text{fwd}}(h_{t}^{e},\overrightarrow{h_{t-1}})$ (3)
$\overleftarrow{h_{t}}=\text{LSTM}_{\text{bwd}}(h_{t}^{e},\overleftarrow{h_{t+1}})$ (4)
$h_{t}^{\text{enc}}=[\overrightarrow{h_{t}};\overleftarrow{h_{t}}]$ (5)
where $\cdot$$\cdot$] denotes concatenation. The resulting hidden state $h_{t}^{\text{enc}}\in\mathbb{R}^{2\times d_{\text{model}}}$ encapsulates information from both past and future contexts. This is particularly valuable for anomaly detection, as the nature of an event at time $t$ can often be better understood by observing subsequent data points. The module can also be configured with multiple LSTM layers to learn hierarchical temporal features. The output is a sequence of contextualized representations $\mathbf{H}^{\text{enc}}=[h_{1}^{\text{enc}},\dots,h_{T}^{\text{enc}}]^{\top}$.
Anomaly Scoring Module. Finally, the Anomaly Scoring Module maps the contextualized representation $h_{t}^{\text{enc}}$ of each timestamp to a single scalar value. This scalar represents the raw, unnormalized score (logit) indicating the likelihood of that timestamp being anomalous. This is achieved with a simple linear layer (the classifier) applied pointwise to each element of the sequence $\mathbf{H}^{\text{enc}}$:
$s_{t}=\mathbf{W}_{c}h_{t}^{\text{enc}}+b_{c}$ (6)
where $\mathbf{W}_{c}\in\mathbb{R}^{1\times(2\cdot d_{\text{model}})}$ and $b_{c}\in\mathbb{R}$ are the weight and bias of the linear classifier. The model outputs a sequence of anomaly logits $\mathbf{S}=[s_{1},s_{2},\dots,s_{T}]^{\top}\in\mathbb{R}^{T}$. These logits can be converted into probabilities using the sigmoid function, $p_{t}=\sigma(s_{t})$, for interpretation or thresholding.
III-B2 Training Objective
Given the availability of timestamp-level labels $\mathbf{y}=[y_{1},\dots,y_{T}]^{\top}$, we can train the STAND model in an end-to-end supervised fashion. We frame the task as a binary classification problem for each timestamp. The model is optimized by minimizing the Binary Cross-Entropy (BCE) loss between the predicted scores and the ground-truth labels. To improve numerical stability, we introduce a sigmoid layer into the computation of BCE loss. The loss for a single time-series sample $(\mathbf{X},\mathbf{y})$ is calculated as:
$\mathcal{L}(\mathbf{S},\mathbf{y})=-\frac{1}{T}\sum_{t=1}^{T}[y_{t}\cdot\log(\sigma(s_{t}))+(1-y_{t})\cdot\log(1-\sigma(s_{t}))]$ (7)
where $s_{t}$ is the output logit from the model for timestamp $t$, $y_{t}\in\{0,1\}$ is the corresponding ground-truth label, and $\sigma(\cdot)$ is the sigmoid function. The total loss is averaged over all samples in a mini-batch. The model parameters are updated via backpropagation using a standard gradient-based optimizer such as Adam [adam]. The complete training and inference procedures are summarized in Algorithm 1.
Algorithm 1 Training and Inference with STAND
1:Input: Training set $\mathcal{D}_{\text{trn}}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N_{\text{trn}}}$, test sequence $\mathbf{X}_{\text{test}}$, learning rate $\eta$, number of epochs $E$.
2:Output: Anomaly scores $\mathbf{S}_{\text{test}}$ for $\mathbf{X}_{\text{test}}$.
3:
4:procedure Train($\mathcal{D}_{\text{trn}},\eta,E$)
5:  Initialize STAND model parameters $\theta$.
6:  for epoch = 1 to $E$ do
7:   for each batch $(\mathbf{X},\mathbf{y})$ in $\mathcal{D}_{\text{trn}}$ do
8:     # Forward pass
9:     Compute anomaly logits: $\mathbf{S}\leftarrow\text{STAND}(\mathbf{X};\theta)$.
10:     # Compute loss
11:     Compute loss $\mathcal{L}$.
12:     # Backward pass and optimization
13:     Update parameters: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$.
14:   end for
15:  end for
16:  return Trained parameters $\theta$.
17:end procedure
18:
19:procedure Inference($\mathbf{X}_{\text{test}},\theta$)
20:  Load STAND model with trained parameters $\theta$.
21:  Compute anomaly logits: $\mathbf{S}_{\text{test}}\leftarrow\text{STAND}(\mathbf{X}_{\text{test}};\theta)$.
22:  return Anomaly scores $\mathbf{S}_{\text{test}}$.
23:end procedure
IV Theoretical Analysis
To theoretically substantiate the efficacy of the proposed STAND framework, we provide a rigorous examination of its computational efficiency and optimization stability. This analysis focuses on the asymptotic complexity relative to input dimensions and establishes the convergence properties of the optimization objective.
IV-A Computational Complexity Analysis
The computational efficiency of STAND is a direct consequence of its streamlined architecture, which eschews the quadratic complexity often associated with canonical attention mechanisms.
Let the input time series be denoted by $\mathbf{X}\in\mathbb{R}^{T\times C}$, where $T$ represents the temporal length and $C$ is the number of variables. We denote the hidden dimension of the model as $d$ and the number of layers in the Temporal Encoding Module as $L$.
The complexity of the Feature Embedding module, which projects inputs via a multi-layer perceptron, is computed pointwise. For the entire sequence, the computational cost $\mathcal{C}_{\text{embed}}$ is formulated as:
$\mathcal{C}_{\text{embed}}=\mathcal{O}(T\cdot C\cdot d)$ (8)
Following the embedding, the bidirectional LSTM employed in the Temporal Encoding Module processes the sequence recurrently. The state update for a single LSTM cell involves matrix-vector multiplications proportional to the square of the hidden dimension. Consequently, for a sequence of length $T$ and a bidirectional configuration across $L$ layers, the temporal processing cost $\mathcal{C}_{\text{temp}}$ is bounded by:
$\mathcal{C}_{\text{temp}}=\mathcal{O}(T\cdot d^{2}\cdot L)$ (9)
The Anomaly Scoring Module acts as a linear projection from the hidden space to the scalar anomaly score, contributing a cost $\mathcal{C}_{\text{score}}$:
$\mathcal{C}_{\text{score}}=\mathcal{O}(T\cdot d)$ (10)
By aggregating these components, the total time complexity per training epoch, denoted as $\mathcal{T}_{\text{total}}$, is derived as:
$\mathcal{T}_{\text{total}}=\mathcal{C}_{\text{embed}}+\mathcal{C}_{\text{temp}}+\mathcal{C}_{\text{score}}\approx\mathcal{O}(T(Cd+Ld^{2}))$ (11)
This formulation highlights a critical theoretical advantage: the complexity scales linearly with the sequence length $T$. In stark contrast, state-of-the-art unsupervised methods relying on self-attention mechanisms, such as Anomaly Transformer, exhibit a quadratic complexity $\mathcal{T}_{\text{Trans}}$:
$\mathcal{T}_{\text{Trans}}\approx\mathcal{O}(T^{2}\cdot d)$ (12)
This makes $\mathcal{T}_{\text{Trans}}\gg\mathcal{T}_{\text{total}}$ for long sequence inputs ($T\gg d$).
Regarding memory constraints, the space complexity consists of parameter storage $\mathcal{M}_{\text{param}}$ and runtime activation memory $\mathcal{M}_{\text{run}}$. The parameter space is dominated by the LSTM weights:
$\mathcal{M}_{\text{param}}=\mathcal{O}(L\cdot d^{2}+C\cdot d)$ (13)
IV-B Convergence Analysis
Beyond efficiency, the utilization of supervision signals introduces a stable gradient flow that facilitates convergence. We formulate the learning objective as an Empirical Risk Minimization (ERM) problem.
Let $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ be the training dataset. The objective function $J(\theta)$ is defined as the average Binary Cross-Entropy loss over the dataset:
$J(\theta)=-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left[y_{i,t}\log(\sigma(s_{i,t}))+(1-y_{i,t})\log(1-\sigma(s_{i,t}))\right]$ (14)
where $\theta$ represents the set of learnable parameters, and $s_{i,t}$ is the logit output at time $t$.
To analyze convergence, we assume that the gradient of the objective function, $\nabla J(\theta)$, is $\mathcal{K}$-Lipschitz continuous. This implies that for any two parameter vectors $\theta_{1},\theta_{2}$, the following inequality holds:
$\|\nabla J(\theta_{1})-\nabla J(\theta_{2})\|\leq\mathcal{K}\|\theta_{1}-\theta_{2}\|$ (15)
where $\mathcal{K}>0$ is the Lipschitz constant.
Consider the parameter update rule using standard Gradient Descent with a learning rate $\eta$:
$\theta_{k+1}=\theta_{k}-\eta\nabla J(\theta_{k})$ (16)
According to the Descent Lemma for Lipschitz continuous functions, the objective value at the next iteration is bounded by:
$J(\theta_{k+1})\leq J(\theta_{k})+\langle\nabla J(\theta_{k}),\theta_{k+1}-\theta_{k}\rangle+\frac{\mathcal{K}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$ (17)
Substituting the update rule into this inequality yields:
$J(\theta_{k+1})$ $\leq J(\theta_{k})-\eta\|\nabla J(\theta_{k})\|^{2}+\frac{\mathcal{K}\eta^{2}}{2}\|\nabla J(\theta_{k})\|^{2}$ (18)
$=J(\theta_{k})-\eta\left(1-\frac{\mathcal{K}\eta}{2}\right)\|\nabla J(\theta_{k})\|^{2}$
Assuming the learning rate is chosen such that $\eta<\frac{2}{\mathcal{K}}$, the term $\eta(1-\frac{\mathcal{K}\eta}{2})$ is strictly positive. This guarantees that the sequence of loss values $\{J(\theta_{k})\}$ is non-increasing.
By summing the inequality over $K$ iterations and rearranging terms, we obtain:
$\sum_{k=0}^{K}\eta\left(1-\frac{\mathcal{K}\eta}{2}\right)\|\nabla J(\theta_{k})\|^{2}\leq J(\theta_{0})-J(\theta_{K+1})$ (19)
Since the Binary Cross-Entropy loss is bounded below by 0, i.e., $J(\theta)\geq 0$, the right-hand side is bounded by $J(\theta_{0})$. Taking the limit as $K\to\infty$, the series converges:
$\lim_{K\to\infty}\sum_{k=0}^{K}\|\nabla J(\theta_{k})\|^{2}<\infty$ (20)
This implies that the gradient norm must approach zero:
$\lim_{k\to\infty}\|\nabla J(\theta_{k})\|=0$ (21)
While global optimality is not theoretically guaranteed due to the non-convex nature of the recurrent architecture, Eq. (14) rigorously confirms that STAND converges to a critical point (a local minimum or saddle point). The supervised nature of the loss landscape provides steeper gradients around anomalies compared to reconstruction-based unsupervised losses, thereby empirically accelerating this convergence rate.
V Experiments
V-A Task Designs
To validate the motivation of this paper, we establish the following experimental tasks:
  1. 1.
    Comparison of Unsupervised and Supervised Methods: To analyze the performance of the two method categories, we follow the conventional approach and test unsupervised methods under the unsupervised condition. Constrained by the scarcity of labels, we use half of the labeled data as the training set for supervised methods, and test them on the original test set.
  2. 2.
    Analysis of Supervisory Gain: Considering that the previous task might be subject to certain fairness concerns (due to using 50% labeled data), we further reduce the amount of labeled time series used for training (utilizing as little as 10% of the data) and test on the original test set. We use four subsets of the PSM dataset (SP1-SP4) with increasing anomaly counts (from least to most) for this analysis. The specific dataset partitioning method and data characteristics are introduced in Sec. V-B3.
  3. 3.
    Effectiveness Analysis of STAND: To verify the effectiveness of our proposed STAND method, we conduct ablation experiments on its key components.
  4. 4.
    Sensitivity Analysis of STAND: To explore the influence of parameters on STAND, we perform an analysis of important hyperparameters.
  5. 5.
    Visualization Analysis: To further analyze the advantages of supervised methods, we conduct a visualization analysis of the outputs from different models.
V-B Task Setups
V-B1 Metrics
To evaluate the global anomaly detection capability of the different models, we adopt the point-wise evaluation metrics, F1 score and Area Under the Receiver Operating Characteristic Curve (AUC-ROC).
To assess the models’ event-level anomaly detection capability, we introduce the Aff-F1, UAff-F1 [simad], and Volume Under the Surface for the Precision-Recall surface (VUS-PR) [VUS] scores. Aff-F1 evaluates the interval-level F1-score by utilizing interval membership to quantify the distance between predicted results and ground truth labels. Furthermore, UAff-F1 is an unbiased refinement of Aff-F1, placing greater emphasis on the precision of the model’s detection. VUS-PR is the area under the event-level Precision-Recall curve, which effectively assesses a model’s ability to detect anomalous events.
Furthermore, we introduce Confidence-Consistency Evaluation (CCE) to quantify the prediction consistency of the models [cce]. This metric not only considers the accuracy of the anomaly scores but also accounts for local and global prediction uncertainty. It has been theoretically proven to be more robust and is particularly suitable for scenarios where anomaly labels or anomaly scores are noisy.
V-B2 Baselines
The UTAD-I methods consist of traditional unsupervised anomaly detection approaches, covering Random Guess (Random), a method that predicts each time point as anomalous with a 50% probability and serves as an intuitive baseline to contrast the performance gaps between other methods and a random strategy, Isolation Forest (IForest), Local Outlier Factor (LOF), Principal Component Analysis (PCA), Histogram-based Outlier Score (HBOS), K-Nearest Neighbors (KNN), and K-Means Clustering (KMeans). These methods rely on classical statistical characteristics, distance metrics, or clustering mechanisms to identify anomalies without requiring labeled training data.
The UTAD-II methods encompass a broader range of techniques, including traditional kernel methods, deep learning models, and Transformer-based approaches, specifically including One-Class Support Vector Machine (OCSVM), Autoencoder (AE), Convolutional Neural Network (CNN), Long Short-Term Memory Network (LSTM), TranAD [trans_ad], USAD [usad], Omni [omni], Anomaly Transformer (A.T.) [anom_trans], TimesNet [timesnet], M2N2 [M2N2], LFTSAD [LFTSAD], and CATCH [catch]. This category integrates shallow kernel-based learning, deep neural networks (for spatial-temporal feature extraction), and state-of-the-art Transformer architectures (for capturing long-range dependencies in time series).
The STAD methods are composed of traditional supervised learning classification models, including Random Forest (RF), Support Vector Machine (SVM), AdaBoost, Extra-Trees Classifier (ExtraTrees), and Histogram-based Gradient Boosting Classification Tree (LightGBM). These models are trained on labeled normal/anomalous samples to learn discriminative patterns for direct anomaly classification.
Implementation Details:
  • All methods in UTAD-I, as well as OCSVM, AE, CNN, LSTM, TranAD, USAD, Omni, Anomaly Transformer (A.T.), and TimesNet in UTAD-II, adopt the implementations provided by the TSB-AD benchmark library [^2] https://github.com/TheDatumOrg/TSB-AD. TSB-AD is a widely recognized time-series anomaly detection benchmark that ensures consistent experimental settings (e.g., hyperparameter ranges, data preprocessing) across different models, enhancing the reproducibility of our comparisons.
  • For M2N2, LFTSAD, and CATCH in UTAD-II, we implement the models based on the official open-source code released by the authors of their original papers. We adhere to the recommended hyperparameter configurations, network architectures, and training protocols reported in the literature to maintain the authenticity of the baseline performances.
  • All methods in STAD are implemented based on the Scikit-learn library [^3] https://scikit-learn.org/, a mature and extensively validated machine learning toolkit in the Python ecosystem. The Scikit-learn implementations are chosen for their stability, efficiency, and widespread adoption in academic and industrial research, ensuring reliable baseline results.
V-B3 Datasets
To evaluate the performance of different algorithms in real-world scenarios, we utilize five multivariate time series anomaly detection datasets: PSM, SWaT, WADI, Swan, and Water [dcdetector, simad]. The characteristics of these datasets are summarized in Table I.
For Task 2, to further investigate the gain conferred by supervision, we select PSM as the base dataset. We partition the original test set’s initial portion to serve as the training set, aiming to verify the impact of supervisory labels on the model. This process results in the four PSM subsets shown in Table II (i.e., SP1, SP2, SP3, and SP4). It is important to note that only PSM satisfies the requirement that anomalies are present throughout the entire test set. The anomaly distribution patterns in the other datasets lack sufficient regularity, making it difficult to partition the dataset—specifically, splitting the entire test set into two continuous segments. Hence, we exclusively chose PSM as the base dataset for this task.
Specifically, we perform the dataset partitioning operation as follows: We first select a specific Split Threshold, such as $10\%$ or $20\%$, which indicates that the anomaly label ratio in the training set must be at least greater than this threshold. For instance, SP1 selects $10\%$ as the split threshold. Its resulting training set has an anomaly ratio of $13.35\%$ and contains $79,057$ remaining time steps of data.
TABLE I: Characteristics of the five real-world datasets. #Channel denotes the number of variables in the time series data. #Train and #Test represent the number of time steps in the training and test sets, respectively.
Dataset #Channel #Train #Test Anomaly Rate
PSM 25 132481 87841 27.76%
SWaT 50 495000 449919 12.14%
WADI 122 241921 34561 5.74%
Swan 38 60000 60000 32.60%
Water 9 69260 69261 1.05%
TABLE II: Characteristics of the PSM subsets.
Dataset Split Threshold #Train Training Anomaly Rate #Remaining Remaining Anomaly Rate #Overall Overall Anomaly Rate
SP1 10.00% 8784 13.35% 79057 29.36% 87841 27.76%
SP2 20.00% 20182 20.00% 67659 30.07%
SP3 30.00% 26352 38.73% 61489 23.05%
SP4 40.00% 35136 30.73% 52705 25.77%
V-C Comparison of Unsupervised and Supervised Methods
Table III presents the results for a selection of the better-performing UTAD-I, UTAD-II, and STAD methods; the full results are available in Appendix A. It is evident that unsupervised methods (UTAD-I and UTAD-II) are rarely the top-performing approaches across all datasets. Furthermore, they often exhibit suboptimal overall performance on specific metric categories, which suggests limitations in their practical utility. For instance, recent methods like M2N2 and CATCH show low CCE scores across all five datasets, while methods such as Omni and A.T. achieve low Aff-F1 and UAff-F1 scores on PSM, SWaT, and Swan.
Figure 2 illustrates the average performance of different method categories across multiple datasets and metrics. Although UTAD-II, which incorporates the “normal-only” prior, achieves slightly better average results than UTAD-I, their average performance is not even competitive with simple baselines like PCA.
In contrast to UTAD-I and UTAD-II, STAD methods show a substantial improvement in overall integrated performance, elevating the score by $28.83$ and $24.02$ respectively. Among all methods, our proposed STAND baseline demonstrates the best comprehensive performance. This strongly suggests that STAD methods possess significantly greater potential compared to UTAD approaches. Additionally, the average CCE score for all UTAD models does not exceed $10$, which is far lower than that of the STAD category. This indicates poor prediction consistency for UTAD methods, leading to insufficient practical availability in real-world scenarios.
This comparison experiment was initially designed to maintain consistency with previous related work. However, conducting it under supervised conditions might introduce an element of unfair comparison. Therefore, we further adopt a fairer setting (i.e., evaluating models only on the entirely unseen test set) to compare UTAD and STAD methods in Appendix B. The analysis shows that regardless of the experimental setting, the overall performance of STAD is superior to both UTAD-I and UTAD-II, and the performance of STAND significantly surpasses the current state-of-the-art methods.
Finally, STAND exhibits relatively stronger performance on SWaT and WADI (datasets with larger data volume) compared to other STAD models, but performs less effectively on smaller datasets. This observation suggests that when the data volume is small but a limited number of labels are available, classical STAD methods are a better choice; however, when the data volume is large, the deep learning-based STAND can achieve superior results.
TABLE III: Performance comparison of the three method categories across five datasets and six metrics. Bold indicates the best performance, and underlining indicates the second-best performance. Methods highlighted in gray are UTAD-I, methods highlighted in green are UTAD-II, and methods highlighted in orange represent STAD methods.
Dataset PSM SWaT WADI
Metric CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR
Random -0.83 8.51 67.99 5.53 50.10 32.99 -0.48 7.07 68.96 -3.13 49.96 12.77 -0.07 4.90 69.16 -6.51 49.80 8.47
IForest -5.40 3.75 49.60 -1.69 46.76 29.96 -2.02 6.02 62.11 -9.86 32.74 10.37 7.83 10.99 61.99 -11.06 74.70 17.17
PCA 6.56 18.77 57.93 52.38 74.13 55.01 4.52 48.84 66.17 30.80 89.73 61.73 14.39 39.07 81.74 58.92 81.07 37.04
HBOS -3.44 8.52 44.79 -18.88 49.90 33.23 9.65 10.38 71.49 24.54 75.08 26.34 9.44 9.86 67.40 6.38 73.97 18.50
KMeans 0.12 7.85 60.05 44.55 38.64 30.14 4.79 19.49 78.12 46.24 33.06 18.18 6.46 12.72 66.56 28.87 64.70 13.79
AE 7.78 5.19 56.50 51.07 56.50 34.90 16.25 17.87 74.95 40.34 84.69 45.08 15.04 27.00 74.26 42.28 64.38 16.65
CNN 0.57 26.33 69.16 60.48 61.26 52.52 29.04 58.34 5.46 5.46 86.77 60.84 2.39 24.20 78.57 43.04 66.08 24.00
LSTM 0.60 26.86 66.48 57.44 63.88 51.85 29.54 58.32 5.41 5.41 86.63 68.63 2.47 23.98 78.70 43.07 70.80 25.67
TranAD 0.23 20.09 63.50 58.40 60.54 46.85 29.78 58.34 5.34 5.34 88.47 63.50 2.60 23.23 77.80 40.01 72.76 30.27
USAD 1.49 21.39 46.14 44.90 61.53 46.62 29.79 58.34 5.33 5.33 88.71 63.04 8.36 23.01 77.83 39.86 73.60 28.02
Omni 3.93 21.78 23.44 23.04 64.11 44.55 29.80 58.34 5.31 5.31 88.85 59.30 10.84 39.02 72.07 52.59 79.13 38.82
A.T. 1.02 18.20 48.73 42.56 59.13 43.46 29.31 58.34 5.35 5.35 74.39 47.13 3.88 28.02 74.06 22.00 61.97 20.27
M2N2 -0.07 15.26 79.11 47.11 63.35 48.17 0.01 23.99 75.65 42.15 81.15 33.29 1.69 20.75 74.89 22.97 64.38 22.25
CATCH 0.62 16.85 84.85 62.92 60.65 51.99 0.02 9.33 73.41 23.84 29.78 11.87 0.26 13.64 74.92 31.67 69.76 21.18
RF 44.33 64.86 76.49 69.43 90.77 80.86 61.42 56.73 86.06 85.08 95.92 89.24 61.83 87.71 90.38 88.91 97.58 91.77
SVM 15.66 29.80 22.34 21.81 91.93 80.46 26.95 35.33 81.49 80.25 93.29 88.62 32.21 80.41 92.49 84.41 93.89 87.95
AdaBoost 14.46 30.48 29.28 28.95 84.99 75.80 18.69 52.01 50.13 42.59 86.71 41.43 24.53 78.40 83.81 78.44 91.84 75.43
ExtraTrees 50.28 66.16 78.92 73.20 92.82 83.89 69.03 53.87 81.26 81.22 97.01 92.69 63.07 84.94 87.16 79.12 99.10 95.48
LightGBM 44.21 30.53 12.18 12.18 87.36 77.24 66.51 48.82 75.89 70.81 94.87 80.76 66.43 80.95 91.93 89.76 96.89 92.03
STAND 46.94 30.33 29.65 29.35 90.74 83.66 71.48 57.95 87.84 87.42 92.02 87.14 65.24 82.03 89.93 84.07 91.72 87.08
Dataset Swan Water Average
Metric CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR
Random -0.27 8.65 26.68 -0.69 49.79 83.90 0.72 1.53 63.75 -3.53 50.32 3.57 -0.18 6.13 59.31 -1.67 49.99 28.34
IForest 5.85 18.18 11.56 5.15 66.30 85.11 12.35 8.11 60.59 19.21 79.10 10.09 3.72 9.41 49.17 0.35 59.92 30.54
PCA 1.55 24.75 2.21 2.00 59.62 93.45 11.38 18.03 62.58 47.28 90.68 17.92 7.68 29.89 54.12 38.28 79.05 53.03
HBOS 15.92 18.01 23.08 20.14 78.82 88.77 16.45 8.16 65.15 9.51 82.93 11.87 9.60 10.99 54.38 8.34 72.14 35.74
KMeans -7.41 6.43 11.18 6.27 30.04 81.24 28.38 17.89 72.45 46.33 92.85 24.01 6.47 12.88 57.67 34.45 51.86 33.47
AE 5.95 18.91 21.89 20.80 60.01 88.57 1.76 5.68 54.44 22.81 50.71 2.36 9.36 14.93 56.41 35.46 63.26 37.51
CNN 0.56 24.84 15.92 15.78 72.99 93.71 2.62 3.96 44.78 9.19 54.04 2.55 7.04 27.53 42.78 26.79 68.23 46.72
LSTM 0.38 23.98 27.41 27.11 75.24 92.63 2.62 8.63 54.09 12.65 60.35 5.84 7.12 28.36 46.42 29.14 71.38 48.92
TranAD 0.08 15.94 18.57 5.76 65.01 91.96 0.32 6.20 50.82 13.96 52.11 6.51 6.60 24.76 43.21 24.69 67.78 47.82
USAD 0.54 20.85 11.95 11.42 61.81 92.27 3.21 9.35 41.21 13.84 71.66 6.27 8.68 26.59 36.49 23.07 71.46 47.25
Omni 1.86 24.44 0.26 -0.25 68.55 94.25 3.23 9.30 41.19 5.74 74.37 6.44 9.93 30.58 28.45 17.29 75.00 48.67
A.T. 0.02 3.45 6.64 -5.31 50.27 84.22 1.91 3.29 71.44 30.10 65.94 5.02 7.23 22.26 41.24 18.94 62.34 40.02
M2N2 0.42 26.50 0.03 0.03 77.54 96.62 0.65 22.13 84.87 63.64 91.91 37.77 0.54 21.73 62.91 35.18 75.67 47.62
CATCH 0.22 25.60 3.05 3.03 61.55 94.17 0.88 18.98 81.54 54.51 91.28 27.36 0.40 16.88 63.55 35.19 62.61 41.31
RF 62.27 45.36 81.28 80.84 96.39 97.94 70.84 39.31 81.52 78.35 93.55 55.41 50.75 37.69 71.85 66.12 92.29 75.20
SVM 21.99 25.79 51.55 51.29 89.07 97.04 51.56 32.48 98.42 96.81 98.55 82.80 21.22 28.20 61.82 53.88 86.47 67.16
AdaBoost 27.03 25.56 41.22 41.03 80.13 92.86 47.29 32.58 92.36 82.25 98.83 88.42 23.21 29.25 53.55 44.76 82.25 61.13
ExtraTrees 57.48 41.98 74.27 73.10 95.58 97.51 92.52 35.80 98.42 96.62 99.76 93.93 50.19 34.57 69.58 60.87 91.47 73.90
LightGBM 67.56 25.14 46.38 46.04 94.28 97.34 71.07 73.33 82.77 79.06 94.11 54.01 47.94 32.42 63.68 51.93 88.72 68.99
STAND 55.53 25.57 46.98 46.65 84.76 97.19 84.44 34.01 96.61 93.41 99.62 95.38 64.73 45.98 70.20 68.18 91.77 90.09
V-D Analysis of Supervisory Gain
To explicitly demonstrate the impact of supervisory information on model performance, we train models using four subsets of the PSM dataset (SP1-SP4) and evaluate them on the original, full PSM test set. Notably, SP1 contains only $10\%$ of the PSM data. The performance of different models on these four subsets is summarized in Table IV.
Strikingly, the average performance of ExtraTrees, when trained on the SP1 subset ($10\%$ of the data), reaches $44.19$, which already surpasses the best UTAD-I method, PCA (average score $44.13$). Moreover, it maintains considerable competitiveness against the best UTAD-II method, CATCH. When $20\%$ of the data (i.e., SP2) is used for training, the average performance of ExtraTrees improves to $53.03$, representing an increase of $14.5\%$ compared to CATCH. Furthermore, when $40\%$ of the data (i.e., SP4) is utilized for training, ExtraTrees shows an average improvement of $62.66\%$ relative to CATCH.
These results provide compelling evidence of the massive gain conferred by supervisory information. They clearly demonstrate that fully exploiting label information yields greater performance benefits than merely pursuing complex model architecture improvements.
TABLE IV: Performance of all supervised methods across the four PSM subsets (SP1–SP4).
Model Source CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR Avg.
PCA PSM 6.56 18.77 57.93 52.38 74.13 55.01 44.13
CATCH PSM 0.62 16.85 84.85 62.92 60.65 51.99 46.31
RF SP1 13.73 15.60 56.82 49.29 66.79 45.75 41.33
SVM 8.28 15.75 45.52 38.47 68.93 47.72 37.45
AdaBoost 6.39 20.37 40.49 36.55 67.03 46.54 36.23
ExtraTrees 17.06 21.56 56.08 47.21 71.68 51.54 44.19
LightGBM 15.63 15.79 63.16 41.29 71.23 50.37 42.91
STAND 17.84 25.16 52.05 35.07 70.46 54.69 42.54
RF SP2 28.75 30.31 25.14 24.83 82.24 64.35 42.60
SVM 15.33 29.60 25.69 21.27 84.38 65.26 40.26
AdaBoost 10.01 18.40 25.37 21.89 76.14 49.52 33.55
ExtraTrees 33.31 30.34 48.76 48.76 85.84 71.13 53.03
LightGBM 32.35 30.38 26.44 25.44 84.74 68.87 44.70
STAND 29.92 27.56 44.93 40.48 84.83 71.37 49.85
RF SP3 33.83 56.66 30.91 29.21 84.85 65.61 50.18
SVM 14.46 4.65 25.62 18.01 83.02 59.35 34.18
AdaBoost 10.78 16.70 27.22 21.90 74.58 48.81 33.33
ExtraTrees 36.45 57.12 53.36 52.30 85.47 66.97 58.61
LightGBM 33.02 22.55 6.82 6.36 81.85 60.27 35.14
STAND 34.44 26.58 42.81 39.48 87.48 69.41 50.03
RF SP4 44.26 63.97 74.74 67.03 90.62 81.21 70.30
SVM 15.66 29.80 22.34 21.79 91.93 80.47 43.66
AdaBoost 11.30 16.11 27.28 22.58 79.42 54.10 35.13
ExtraTrees 50.25 66.67 81.67 76.72 92.56 84.11 75.33
LightGBM 45.30 30.53 12.57 12.57 89.30 78.37 44.77
STAND 42.67 30.46 28.41 28.31 91.14 82.36 50.56
V-E Effectiveness Analysis of STAND
To analyze the effectiveness of the bidirectional hidden state (Bidir.) and the Temporal Encoding Module (TEM) within STAND for anomaly detection, we conducted ablation experiments on these two components. The results are presented in Table V.
When the TEM is removed, the model’s performance significantly degrades across almost all datasets and metrics, with the average score dropping by $37.2\%$. This result underscores that the TEM is the most crucial component of the STAND framework. If the TEM is retained but the Bidir. is disabled, the overall performance decreases by $8.07\%$. This indicates that multi-directional modeling provides additional effective information for the anomaly detection task.
TABLE V: Ablation results of STAND on five datasets. Bidir. with indicates that only the forward hidden state in Eq. (5) is used. TEM with indicates that the Temporal Encoding Module is replaced by an Identity function. Bold and underlining denote the best and second-best performing models, respectively, within a single dataset.
Bidir. TEM Dataset CCE F1 Aff-F1 UAff-F1 AUC V-PR
PSM 9.25 21.08 68.19 42.53 67.37 52.51
38.73 29.93 29.13 28.57 88.44 79.77
46.94 30.33 29.65 29.35 90.74 83.66
SWaT 20.04 43.13 84.81 60.83 76.32 51.76
51.95 58.01 89.12 88.59 90.27 80.63
71.48 57.95 87.84 87.42 92.02 87.14
WADI 14.82 30.07 82.49 53.15 76.65 33.51
60.31 76.48 84.22 74.70 95.07 85.94
65.24 82.03 89.93 84.07 91.72 87.08
Swan 3.14 13.35 32.26 24.52 53.58 88.27
51.08 25.37 33.81 33.54 82.56 97.06
55.53 25.57 46.98 46.65 84.76 97.19
Water 8.92 6.43 88.54 76.25 63.13 6.27
64.04 31.62 83.89 64.44 98.64 84.92
84.44 34.01 96.61 93.41 99.62 95.38
Average 11.23 22.81 71.26 51.46 67.41 46.46
53.22 44.28 64.04 57.97 91.00 85.66
64.73 45.98 70.20 68.18 91.77 90.09
Figure 4: Sensitivity analysis of model performance with respect to three key hyperparameters on five datasets. (a)-(b) show the impact of model dimension ($d_{\text{model}}$) on CCE and VUS-PR, respectively; (c)-(d) illustrate the effect of the number of TEM layers; (e)-(f) depict the influence of window size.
Figure 5: Sensitivity analysis of model performance with respect to three key hyperparameters on four subsets of the PSM dataset. (a)-(b) show the impact of model dimension ($d_{\text{model}}$) on CCE and VUS-PR, respectively; (c)-(d) illustrate the effect of the number of TEM layers; (e)-(f) depict the influence of window size.
V-F Sensitivity Analysis of STAND
To investigate the influence of hyperparameters on STAND, we search for the model’s CCE and VUS-PR performance across various parameter settings on all datasets and the PSM subsets. The results are presented in Figure 4 and Figure 5, respectively, where the error bars represent the 95% confidence interval (all metric results are provided in Appendix C). Overall, the parameter sensitivity of STAND remains largely consistent, regardless of the dataset or the anomaly ratio within the dataset.
Based on Figure 4(a-b) and Figure 5(a-b), increasing the model dimension ($d_{\text{model}}$) effectively enhances the model’s detection consistency (CCE) and local detection capability (VUS-PR) across different datasets.
From Figure 4(c-d) and Figure 5(c-d), it can be observed that increasing the number of model layers is only effective when the data volume is relatively small (e.g., Swan, SP1). For most scenarios, selecting a single layer or a two-layer model yields better results. Similarly, as shown in Figure 4(e-f) and Figure 5(e-f), a window size of $32$ performs distinctly better than a window size of $128$.
Therefore, overall, learning the relationships between different channels in a high-dimensional space is more effective than deepening the model or enlarging the temporal window size.
V-G Visualization Analysis
Figure 6: The performance of competitive models on the entire SWaT test dataset.
Figure 7: Visualization results for the initial segment (time steps $150,000$ to $200,000$) of the partitioned SWaT dataset.
Figure 8: Visualization results for the latter segment (time steps $300,000$ to $350,000$) of the partitioned SWaT dataset.
To further analyze the advantages of STAD methods in terms of practical utility in real-world scenarios, we selected and visualized the outputs of two better-performing UTAD-II models (LSTM and M2N2) along with three STAD models (ExtraTrees, LightGBM, and STAND) on the SWaT dataset.
In Figure 6, the first two rows represent the time series data for two selected channels, while the subsequent rows show the anomaly scores output by the different models. Red highlighting indicates the ground truth anomaly regions, and the red dashed line signifies the division point for the STAD training set.
Figures 7 and 8 are localized visualizations of Figure 6, depicting two key regions corresponding to the left and right sides of the red dashed line, respectively. The meaning of each subplot within these figures is consistent with that of Figure 6.
Overall, the average performance scores for LSTM and M2N2 on SWaT are $68.63$ and $33.29$, respectively. However, in practice, only the long-duration anomaly segment around time step $250,000$ is predicted well. Anomalies in other locations are almost entirely missed. In contrast, all three STAD models successfully predict the majority of anomalies within the training set portion. Crucially, they are also capable of detecting partial anomalies in the unseen test set. Among them, STAND exhibits the lowest rate of false positives, demonstrating superior practical availability. Conversely, the two UTAD-II models fail to predict the anomalies in the latter half of the series.
In the preceding segment (Figure 7), both unsupervised methods (LSTM and M2N2) only output extremely small anomaly scores. This suggests that their anomaly scores are difficult for human interpretation and are prone to causing false negatives (missed detections). Conversely, under the influence of the supervisory signal, the three STAD methods are capable of perfectly predicting the anomalies they have already encountered.
In the subsequent segment (Figure 8), both UTAD methods completely fail to detect the anomalies. Among the STAD approaches, ExtraTrees and STAND successfully detect two anomalous segments, while LightGBM only detects one. However, ExtraTrees and LightGBM also suffer from significant false positives, making it difficult for humans to correctly interpret and locate the true anomalies. In contrast, STAND predicts anomaly segments that are cleaner and more interpretable.
Generally, the UTAD methods did not perform as anticipated; their practical performance on the majority of anomalies is poor. In stark contrast, the outputs of the STAD methods are more reliable and readily interpretable by humans.
VI Conclusion
This paper critically re-evaluates the prevailing research paradigm in time series anomaly detection (TSAD), which heavily favors complex model architectures to address the lack of labels. By proposing STAND, a simple yet effective supervised baseline, we quantify the impact of supervisory signals against architectural sophistication. Our empirical results conclusively show that “labels matter more than models.” Even with a minimal labeling budget (e.g., 10% of data), STAND significantly outperforms existing complex unsupervised methods in terms of detection accuracy, F1 score, and prediction consistency. These findings reveal the limitations of the current algorithm-centric path and highlight the immense value of limited supervision. Consequently, we suggest that future TSAD research should pivot towards data-centric approaches that maximize the utility of available labels to enhance practical deployment performance.
Acknowledgments
This work was supported in part by National Natural Science Foundation of China No. 92467109, U21A20478, National Key R&D Program of China 2023YFA1011601, and the Major Key Project of PCL, China under Grant PCL2025A11 and PCL2025A13.
[Uncaptioned image] Zhijie Zhong received the B.S. degree in 2022 from the Harbin Engineering University, Harbin, China and he is currently pursuing the Ph.D. degree in the School of Future Technology, South China University of Technology, Guangzhou, China. His research interests include data mining, machine learning, time series analysis, anomaly detection, and large language model (LLM).
[Uncaptioned image] Zhiwen Yu (S’06-M’08-SM’14) received the Ph.D. degree from the City University of Hong Kong, Hong Kong, China, in 2008. He is currently a Professor with the School of Computer Science and Engineering, South China University of Technology, China. He has authored or co-authored more than 200 refereed journal articles and international conference papers, including more than 80 articles in the journals of IEEE Transactions, h-index 52, and Google citation 13800. He is an Associate Editor of IEEE Transactions on Systems, Man, and Cybernetics: Systems. He is a Senior Member of ACM and a member of Council of China Computer Federation (CCF).
[Uncaptioned image] Kaixiang Yang (M’21) received the B.S. degree and M.S. degree from the University of Electronic Science and Technology of China and Harbin Institute of Technology, China, in 2012 and 2015, respectively, and the Ph.D. degree from the School of Computer Science and Engineering, South China University of Technology, China, in 2020. He has been a Research Engineer with the 7th Research Institute, China Electronics Technology Group Corporation, Guangzhou, China, from 2015 to 2017, and has been a Postdoctoral Researcher with Zhejiang University from 2020 to 2021. He is now with the School of Computer Science and Engineering, South China University of Technology. His research interests include pattern recognition, machine learning, and industrial data intelligence.
[Uncaptioned image] C. L. Philip Chen (S’88-M’88-SM’94-F’07) received the M.S. degree in electrical engineering from the University of Michigan, Ann Arbor, MI, USA, in 1985, and the Ph.D. degree in electrical engineering from Purdue University, West Lafayette, IN, USA, in 1988. He is the Chair Professor and the Dean of the School of Computer Science and Engineering, South China University of Technology, Guangzhou, China. Being a Program Evaluator of the Accreditation Board of Engineering and Technology Education (ABET) in USA, for computer engineering, electrical engineering, and software engineering programs, he successfully architects the University of Macau’s Engineering and Computer Science programs receiving accreditations from Washington/Seoul Accord through Hong Kong Institute of Engineers (HKIE), which is considered as his utmost contribution in engineering/computer science education for Macau as the former Dean of the Faculty of Science and Technology. His current research interests include cybernetics, systems, and computational intelligence. Dr. Chen is a fellow of AAAS, IAPR, CAA, and HKIE, and a member of the Academia Europaea (AE), the European Academy of Sciences and Arts (EASA), and the International Academy of Systems and Cybernetics Science (IASCYS). He received IEEE Norbert Wiener Award in 2018 for his contribution in systems and cybernetics, and machine learnings. He is also a highly cited researcher by Clarivate Analytics in 2018 and 2019. He was a recipient of the 2016 Outstanding Electrical and Computer Engineers Award from his alma mater, Purdue University, in 1988. He was the Chair of TC 9.1 Economic and Business Systems of International Federation of Automatic Control from 2015 to 2017 and currently is a Vice President of Chinese Association of Automation (CAA). He was the IEEE Systems, Man, and Cybernetics Society President from 2012 to 2013, the Editor-in-Chief for the IEEE TRANSACTIONS ON SYSTEMS, MAN, AND CYBERNETICS: SYSTEMS from 2014 to 2019, and currently, he is the Editor-in-Chief for the IEEE TRANSACTIONS ON CYBERNETICS, and an Associate Editor of IEEE TRANSACTIONS ON ARTIFICIAL INTELLIGENCE and IEEE TRANSACTIONS ON FUZZY SYSTEMS.
Appendix
This is the appendix of Labels Matter More Than Models: Quantifying the Benefit of Supervised Time Series Anomaly Detection.
A Full Results of Unsupervised Methods
Table A6 presents the complete results for all unsupervised methods (including both UTAD-I and UTAD-II categories).
TABLE A6: Performance comparison of all unsupervised methods across five datasets and six metrics. Bold indicates the best performance, and underlining indicates the second-best performance. Methods highlighted in gray are UTAD-I, and methods highlighted in green are UTAD-II.
Dataset PSM SWaT WADI
Metric CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR
Random -0.83 8.51 67.99 5.53 50.10 32.99 -0.48 7.07 68.96 -3.13 49.96 12.77 -0.07 4.90 69.16 -6.51 49.80 8.47
IForest -5.40 3.75 49.60 -1.69 46.76 29.96 -2.02 6.02 62.11 -9.86 32.74 10.37 7.83 10.99 61.99 -11.06 74.70 17.17
LOF 0.39 10.09 73.93 23.52 51.71 35.40 0.06 8.08 72.90 17.03 50.60 13.68 0.93 9.32 74.42 28.97 55.18 11.16
PCA 6.56 18.77 57.93 52.38 74.13 55.01 4.52 48.84 66.17 30.80 89.73 61.73 14.39 39.07 81.74 58.92 81.07 37.04
HBOS -3.44 8.52 44.79 -18.88 49.90 33.23 9.65 10.38 71.49 24.54 75.08 26.34 9.44 9.86 67.40 6.38 73.97 18.50
KNN -1.78 9.08 70.11 31.22 37.39 29.49 -1.95 4.17 70.06 10.90 14.49 7.78 -0.37 5.23 68.68 -6.21 51.76 10.31
KMeans 0.12 7.85 60.05 44.55 38.64 30.14 4.79 19.49 78.12 46.24 33.06 18.18 6.46 12.72 66.56 28.87 64.70 13.79
OCSVM 2.84 17.03 43.42 30.49 52.63 38.46 10.10 4.77 39.06 21.13 77.36 24.78 13.37 32.12 70.40 47.82 81.78 32.00
AE 7.78 5.19 56.50 51.07 56.50 34.90 16.25 17.87 74.95 40.34 84.69 45.08 15.04 27.00 74.26 42.28 64.38 16.65
CNN 0.57 26.33 69.16 60.48 61.26 52.52 29.04 58.34 5.46 5.46 86.77 60.84 2.39 24.20 78.57 43.04 66.08 24.00
LSTM 0.60 26.86 66.48 57.44 63.88 51.85 29.54 58.32 5.41 5.41 86.63 68.63 2.47 23.98 78.70 43.07 70.80 25.67
TranAD 0.23 20.09 63.50 58.40 60.54 46.85 29.78 58.34 5.34 5.34 88.47 63.50 2.60 23.23 77.80 40.01 72.76 30.27
USAD 1.49 21.39 46.14 44.90 61.53 46.62 29.79 58.34 5.33 5.33 88.71 63.04 8.36 23.01 77.83 39.86 73.60 28.02
Omni 3.93 21.78 23.44 23.04 64.11 44.55 29.80 58.34 5.31 5.31 88.85 59.30 10.84 39.02 72.07 52.59 79.13 38.82
A.T. 1.02 18.20 48.73 42.56 59.13 43.46 29.31 58.34 5.35 5.35 74.39 47.13 3.88 28.02 74.06 22.00 61.97 20.27
TimesNet 0.01 15.34 73.81 35.76 57.19 46.33 0.12 8.77 74.09 30.50 30.16 11.71 0.59 16.76 73.85 18.57 70.07 20.66
M2N2 -0.07 15.26 79.11 47.11 63.35 48.17 0.01 23.99 75.65 42.15 81.15 33.29 1.69 20.75 74.89 22.97 64.38 22.25
LFTSAD -0.32 9.27 63.57 16.99 44.32 31.62 0.81 19.79 65.12 3.58 66.36 24.21 0.28 9.54 53.74 -27.72 48.41 11.49
CATCH 0.62 16.85 84.85 62.92 60.65 51.99 0.02 9.33 73.41 23.84 29.78 11.87 0.26 13.64 74.92 31.67 69.76 21.18
Dataset Swan Water Average
Metric CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR
Random -0.27 8.65 26.68 -0.69 49.79 83.90 0.72 1.53 63.75 -3.53 50.32 3.57 -0.18 6.13 59.31 -1.67 49.99 28.34
IForest 5.85 18.18 11.56 5.15 66.30 85.11 12.35 8.11 60.59 19.21 79.10 10.09 3.72 9.41 49.17 0.35 59.92 30.54
LOF 0.14 10.32 21.67 5.50 49.14 83.10 0.07 2.91 64.47 1.14 50.30 4.33 0.32 8.14 61.48 15.23 51.39 29.53
PCA 1.55 24.75 2.21 2.00 59.62 93.45 11.38 18.03 62.58 47.28 90.68 17.92 7.68 29.89 54.12 38.28 79.05 53.03
HBOS 15.92 18.01 23.08 20.14 78.82 88.77 16.45 8.16 65.15 9.51 82.93 11.87 9.60 10.99 54.38 8.34 72.14 35.74
KNN -3.17 6.67 20.69 -15.08 32.99 83.35 3.40 5.06 65.87 0.35 55.71 8.86 -0.77 6.04 59.08 4.23 38.47 27.96
KMeans -7.41 6.43 11.18 6.27 30.04 81.24 28.38 17.89 72.45 46.33 92.85 24.01 6.47 12.88 57.67 34.45 51.86 33.47
OCSVM 11.82 10.43 10.96 8.70 71.10 85.26 20.57 17.79 70.93 20.01 92.31 22.78 11.74 16.43 46.95 25.63 75.04 40.66
AE 5.95 18.91 21.89 20.80 60.01 88.57 1.76 5.68 54.44 22.81 50.71 2.36 9.36 14.93 56.41 35.46 63.26 37.51
CNN 0.56 24.84 15.92 15.78 72.99 93.71 2.62 3.96 44.78 9.19 54.04 2.55 7.04 27.53 42.78 26.79 68.23 46.72
LSTM 0.38 23.98 27.41 27.11 75.24 92.63 2.62 8.63 54.09 12.65 60.35 5.84 7.12 28.36 46.42 29.14 71.38 48.92
TranAD 0.08 15.94 18.57 5.76 65.01 91.96 0.32 6.20 50.82 13.96 52.11 6.51 6.60 24.76 43.21 24.69 67.78 47.82
USAD 0.54 20.85 11.95 11.42 61.81 92.27 3.21 9.35 41.21 13.84 71.66 6.27 8.68 26.59 36.49 23.07 71.46 47.25
Omni 1.86 24.44 0.26 -0.25 68.55 94.25 3.23 9.30 41.19 5.74 74.37 6.44 9.93 30.58 28.45 17.29 75.00 48.67
A.T. 0.02 3.45 6.64 -5.31 50.27 84.22 1.91 3.29 71.44 30.10 65.94 5.02 7.23 22.26 41.24 18.94 62.34 40.02
TimesNet 0.40 18.79 12.21 3.41 58.88 92.56 0.25 10.11 63.76 -4.59 71.46 19.50 0.28 13.95 59.54 16.73 57.55 38.15
M2N2 0.42 26.50 0.03 0.03 77.54 96.62 0.65 22.13 84.87 63.64 91.91 37.77 0.54 21.73 62.91 35.18 75.67 47.62
LFTSAD 0.17 10.73 15.32 3.05 51.44 84.95 0.47 4.91 69.24 8.06 45.73 6.68 0.28 10.85 53.40 0.79 51.25 31.79
CATCH 0.22 25.60 3.05 3.03 61.55 94.17 0.88 18.98 81.54 54.51 91.28 27.36 0.40 16.88 63.55 35.19 62.61 41.31
Figure A9: Sensitivity analysis of $d_{\text{model}}$ across 5 datasets on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A10: Sensitivity analysis of number of TEM layers across 5 datasets on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A11: Sensitivity analysis of window size across 5 datasets on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A12: Sensitivity analysis of model dimension $d_{\text{model}}$ across 4 subsets of the PSM dataset on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A13: Sensitivity analysis of number of TEM layers across 4 subsets of the PSM dataset on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A14: Sensitivity analysis of window size across 4 subsets of the PSM dataset on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
B Temporal Split Validation Results
Since the supervised methods utilized the initial segment of the original test set as their training data, it is challenging to fairly analyze whether STAD methods genuinely outperform UTAD methods. To mitigate this drawback, we designate the latter segment of the original test set (after partitioning) as the true evaluation set. Both UTAD and STAD methods are tested on this specific set, and the results are presented in Table A7.
Under this fair comparison setting, the best-performing UTAD-I method is PCA, with an average score of $44.34$. For UTAD-II, the best method is LSTM, achieving an average score of $47.09$. In contrast, recent state-of-the-art UTAD-II methods, M2N2, LFTSAD, and CATCH, yield average scores of $37.91$, $27.96$, and $35.71$, respectively. The two best-performing STAD methods are RF and STAND, with average scores of $45.79$ and $54.68$.
We observe that two simple STAD methods outperform the majority of UTAD methods, and the performance of STAND is significantly superior to both other method categories. Furthermore, we note that recent UTAD-II methods (M2N2, LFTSAD, and CATCH) are surprisingly less competitive than older, simpler approaches (PCA and LSTM). In addition, the prediction consistency (CCE) performance of all UTAD methods is poor, falling far behind that of the STAD approaches. This suggests their limited practical applicability.
TABLE A7: Three method categories’ performance comparison on the latter half of the divided dataset (across five datasets and six metrics). Bold indicates the best performance, and underlining indicates the second-best performance. Methods highlighted in gray are UTAD-I, methods highlighted in green are UTAD-II, and methods highlighted in orange represent STAD methods.
Dataset PSM SWaT WADI
Metric CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR
Random -0.52 8.49 71.43 11.66 50.20 35.30 -0.46 7.87 69.35 -4.59 50.06 19.55 0.99 3.18 66.37 20.87 51.02 6.17
IForest 0.40 14.23 49.22 -1.25 51.34 36.30 0.40 6.72 67.58 -3.93 32.17 15.36 11.57 10.59 72.41 26.20 85.22 19.86
LOF 0.84 11.90 77.47 29.69 52.37 39.09 0.07 7.34 72.47 11.62 49.90 19.73 1.16 7.94 73.75 27.64 54.82 8.02
PCA 10.22 17.77 56.73 51.31 78.32 59.91 5.64 34.75 61.50 -2.99 92.11 77.11 16.32 32.83 83.96 58.69 84.93 29.05
HBOS -0.50 10.66 54.69 8.80 48.40 34.71 14.58 5.07 71.07 26.69 85.23 51.92 16.01 18.53 75.88 35.34 85.02 26.56
KNN -1.87 6.01 70.79 38.02 30.14 26.60 -2.94 1.93 67.76 -2.78 9.41 11.14 -0.12 3.04 68.35 -28.87 57.96 8.82
KMeans -1.10 6.79 63.80 53.36 35.30 28.84 0.53 14.31 77.65 40.88 21.37 17.78 13.50 5.43 60.69 0.36 73.85 10.92
OCSVM -0.60 7.66 52.24 41.95 41.82 31.52 10.49 15.86 40.40 19.28 86.41 45.01 20.36 26.60 70.09 38.04 90.25 30.55
AE 5.94 4.90 35.00 14.38 59.18 37.57 21.36 31.37 76.01 46.24 91.72 76.31 22.08 24.35 73.78 45.87 77.90 18.94
CNN 2.06 24.97 49.02 28.51 78.98 62.25 37.30 41.57 11.57 11.57 91.75 82.35 13.53 28.86 86.69 64.58 87.09 35.37
LSTM 1.87 25.54 53.94 37.36 77.35 61.26 38.07 41.57 12.29 12.29 89.88 81.90 14.65 29.25 86.66 66.06 83.47 35.49
TranAD 0.47 16.40 47.32 43.45 69.66 53.91 38.32 41.57 12.26 12.26 92.69 81.38 12.60 27.93 85.20 66.75 84.60 36.17
USAD 4.28 16.72 44.31 40.91 72.39 55.40 38.30 41.57 12.23 12.23 92.40 80.85 15.29 30.31 86.20 67.15 76.46 29.52
Omni 11.96 18.16 17.55 16.05 72.61 51.57 38.41 41.57 11.16 11.16 92.78 77.46 14.66 30.18 79.33 46.35 78.19 42.14
A.T. 0.51 10.26 63.57 -3.85 50.65 36.48 0.00 1.20 65.24 -14.25 49.26 19.53 -0.34 1.06 35.84 -53.36 36.60 5.76
TimesNet -0.01 12.76 74.16 27.57 55.98 45.34 0.05 4.39 71.47 7.50 19.79 13.50 2.18 15.49 72.32 24.23 70.38 21.66
M2N2 -0.01 14.83 80.84 46.18 65.93 51.71 -0.51 15.32 27.56 25.80 80.08 40.06 10.63 27.13 81.07 38.19 83.14 35.44
LFTSAD 1.02 18.75 63.38 28.61 56.44 44.73 0.73 22.53 70.22 3.36 73.46 45.84 3.30 14.82 69.27 34.03 57.60 13.61
CATCH 0.54 15.26 83.83 56.61 61.46 50.61 0.01 5.79 72.34 10.56 19.74 13.57 2.23 15.75 76.92 39.90 80.45 26.05
RF 13.14 18.81 50.78 21.89 73.17 50.54 31.01 39.04 60.07 56.87 90.84 80.28 23.82 46.73 73.04 60.38 85.49 63.63
SVM 12.25 22.13 65.87 52.31 80.42 57.39 19.93 15.09 68.89 25.96 91.67 76.17 12.87 33.22 81.05 50.59 81.50 42.92
AdaBoost 10.05 20.06 34.14 28.86 70.80 39.94 15.39 22.34 47.91 34.16 84.07 37.79 11.60 27.13 58.41 24.08 78.62 27.07
ExtraTrees 10.78 20.36 53.19 36.77 79.11 57.64 39.10 52.71 12.91 12.89 88.88 53.93 27.43 51.47 88.03 70.48 94.61 73.56
LightGBM 6.79 16.35 55.07 16.70 71.52 51.89 39.25 28.39 30.33 -12.15 89.38 64.47 23.50 44.47 77.52 50.19 86.81 50.02
STAND 16.39 23.58 64.18 55.61 66.81 55.47 51.97 36.88 76.94 69.73 90.68 79.62 26.70 37.33 78.10 55.51 73.17 39.71
Dataset Swan Water Average
Metric CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR
Random -0.10 9.10 26.79 4.24 49.71 89.09 4.74 0.55 66.75 0.38 55.06 0.49 0.93 5.84 60.14 6.51 51.21 30.12
IForest 5.77 17.29 3.23 -3.03 68.63 92.47 37.09 8.07 52.82 -47.52 98.97 17.21 11.05 11.38 49.05 -5.91 67.27 36.24
LOF 0.05 10.84 23.54 14.35 48.62 90.63 6.58 2.10 67.15 0.60 85.45 1.86 1.74 8.02 62.88 16.78 58.23 31.87
PCA 1.52 18.20 1.35 1.33 53.33 96.02 46.62 7.74 77.31 41.62 99.01 37.89 16.06 22.26 56.17 29.99 81.54 60.00
HBOS 13.70 15.04 18.22 13.47 80.02 96.18 31.77 6.85 56.58 -39.18 96.26 9.63 15.11 11.23 55.29 9.02 78.99 43.80
KNN -3.79 7.71 22.44 -15.13 31.37 88.35 22.65 4.20 49.15 -52.68 70.61 5.60 2.79 4.58 55.70 -12.29 39.90 28.10
KMeans -9.32 4.74 16.58 8.88 23.50 87.55 44.24 6.19 39.83 -66.82 93.74 24.05 9.57 7.49 51.71 7.33 49.55 33.83
OCSVM 11.64 11.31 12.23 10.84 68.59 92.06 37.64 8.40 75.02 34.46 99.35 21.63 15.91 13.97 50.00 28.91 77.28 44.15
AE 13.97 13.59 21.21 20.46 68.29 92.43 17.22 1.88 71.53 22.88 58.07 2.88 16.11 15.22 55.51 29.97 71.03 45.63
CNN 0.69 18.42 0.08 0.08 66.83 96.12 31.41 8.51 90.19 78.80 99.87 59.99 17.00 24.47 47.51 36.71 84.90 67.22
LSTM 0.50 16.37 29.49 29.11 71.67 95.29 27.28 5.31 93.18 85.62 91.41 18.73 16.47 23.61 55.11 46.09 82.76 58.53
TranAD 0.08 11.20 19.84 -7.58 60.35 94.03 19.12 5.20 91.42 81.54 94.48 20.85 14.12 20.46 51.21 39.28 80.35 57.27
USAD 0.70 16.44 3.97 3.81 56.80 95.05 29.06 5.53 90.52 79.71 93.95 20.92 17.53 22.11 47.45 40.76 78.40 56.35
Omni 2.45 18.34 0.08 0.08 59.54 95.63 18.24 3.98 92.22 83.31 92.90 42.93 17.14 22.45 40.07 31.39 79.20 61.95
A.T. 0.01 5.64 18.16 -12.74 49.19 89.65 15.94 2.99 51.39 -43.74 61.45 2.99 3.22 4.23 46.84 -25.59 49.43 30.88
TimesNet 0.59 18.48 0.08 0.08 53.38 94.57 0.98 3.76 67.92 10.57 82.65 25.59 0.76 10.97 57.19 13.99 56.44 40.13
M2N2 0.53 18.45 0.08 0.08 85.05 98.41 7.69 5.64 70.36 13.77 91.45 22.53 3.67 16.27 51.98 24.80 81.13 49.63
LFTSAD 0.09 12.80 15.11 -5.85 44.97 90.26 0.11 1.88 51.36 -46.26 51.50 1.12 1.05 14.16 53.87 2.77 56.79 39.11
CATCH 0.40 18.48 0.08 0.08 55.82 95.67 22.06 5.53 75.30 32.23 89.20 44.83 5.05 12.16 61.69 27.88 61.33 46.15
RF 27.51 16.68 29.91 28.55 87.60 95.32 53.80 19.66 65.81 60.20 87.16 26.33 26.14 17.93 54.32 40.44 84.07 51.85
SVM 12.31 16.55 25.65 24.73 82.46 97.19 19.17 7.24 66.59 42.27 65.06 12.07 11.31 12.00 51.72 33.02 77.30 47.20
AdaBoost 33.24 20.19 27.23 25.98 87.31 93.06 33.52 17.43 98.42 96.84 88.12 32.65 17.59 14.29 46.72 28.74 78.36 45.24
ExtraTrees 27.90 17.86 28.26 26.99 87.35 95.54 26.57 13.18 97.91 96.04 85.31 11.49 25.52 18.39 50.63 35.95 84.59 52.82
LightGBM 38.09 16.29 36.29 35.21 89.54 94.90 21.21 26.39 33.03 -2.69 80.66 3.80 24.28 14.79 49.23 21.57 80.89 47.88
STAND 33.26 17.17 27.39 26.95 77.27 96.69 64.19 6.63 92.21 83.33 96.53 20.37 38.50 24.32 67.76 58.22 80.89 58.37
C Full Results of STAND Sensitivity Analysis
To investigate the parameter sensitivity of STAND, we conducted a comprehensive parameter analysis across all five real-world datasets and the four PSM subsets.
Figures A9, A10, and A11 illustrate the impact of the model dimension ($d_{\text{model}}$), the number of TEM layers, and the window size on STAND’s performance across the five datasets, respectively. Overall, the model’s performance improves (for almost all metrics) as the model dimension increases. Conversely, performance tends to degrade (for both point-wise and event-level metrics) as the number of TEM layers and the temporal window size increase.
Figures A9, A10, and A11 also present the impact of $d_{\text{model}}$, the number of TEM layers, and the window size on STAND’s performance across the four PSM subsets. Generally, the trend of model performance change is consistent with that observed on the five real-world datasets. Specifically, when the model dimension increases, the performance improves on three subsets (SP2–SP4), while the performance change on SP1 is relatively marginal. Conversely, when the number of TEM layers and the temporal window size increase, performance slightly decreases on the SP2–SP4 subsets, but no clear trend of performance change is observed on SP1.
Therefore, the influence of hyperparameters on STAND is primarily manifested as follows: When the dataset size is large, the impact is significant, exhibiting a clear performance trend; when the data volume is small, the model is more severely affected by random initialization, leading to less discernible performance trends.
Furthermore, it is important to note that the average level of Aff-F1 and UAff-F1 scores is noticeably better when the training set is SP1 compared to SP2–SP4. This is because these two metrics require the setting of an anomaly evaluation buffer, where predictions within the buffer zone are also regarded as correct. When the anomaly ratio is low, this type of tolerance metric can easily lead to inaccurate evaluations.