### v2

# Labels Matter More Than Models: Rethinking the Unsupervised Paradigm in Time Series Anomaly Detection

Abstract
Time series anomaly detection (TSAD) is a critical data mining task often constrained by label scarcity. Consequently, current research predominantly focuses on Unsupervised Time-series Anomaly Detection (UTAD), relying on increasingly complex architectures to model normal data distributions. However, this algorithm-centric trend often overlooks the significant performance gains achievable from limited anomaly labels available in practical scenarios. This paper challenges the premise that algorithmic complexity is the optimal path for TSAD. Instead of proposing another intricate unsupervised model, we present a comprehensive benchmark and empirical study to rigorously compare supervised and unsupervised paradigms. To isolate the value of labels, we introduce STAND, a deliberately minimalist supervised baseline. Extensive experiments on five public datasets demonstrate that: (1) Labels matter more than models: under a limited labeling budget, simple supervised models significantly outperform complex state-of-the-art unsupervised methods; (2) Supervision yields higher returns: the performance gain from minimal supervision far exceeds the incremental gains from architectural innovations; and (3) Practicality: STAND exhibits superior prediction consistency and anomaly localization compared to unsupervised counterparts. These findings advocate for a paradigm shift in TSAD research, urging the community to prioritize data-centric label utilization over purely algorithmic complexity. The code and benchmark are publicly available at https://github.com/EmorZz1G/STAND.
I Introduction
In the era of big data, the proliferation of Internet of Things (IoT) devices and industrial sensors has generated massive volumes of streaming data [8, 22, 13, 10]. Consequently, Time Series Anomaly Detection (TSAD) has emerged as a crucial and challenging task in big data analytics, with broad applications in industrial system monitoring, cybersecurity, and health surveillance [50, 49, 51, 28, 46].
As the volume and velocity of time series data explode, acquiring high-quality anomaly labels becomes prohibitively expensive. Due to this severe label scarcity within massive datasets, unsupervised time series anomaly detection (UTAD) methods have garnered significant attention in recent years [48, 51, 41, 16]. Typically, unsupervised methods assume that the training time series data primarily consists of normal samples. They detect anomalies by learning the patterns of these normal samples, and this approach currently represents the mainstream of research. Because these methods introduce the prior assumption that the training data is composed entirely of normal samples, they are also referred to as semi-supervised techniques in some literature [28, 41, 26]. Moving forward, we will denote this specific category as Unsupervised Time series Anomaly Detection Type II (UTAD-II), as illustrated in Figure 1(b).
Conversely, earlier studies trained and detected anomalies using unlabeled datasets that might contain anomaly samples, leading to their original designation as unsupervised techniques [28, 31]. We will refer to this category as Unsupervised Time series Anomaly Detection Type I (UTAD-I), as shown in Figure 1(a).
Although there may be some debate regarding the precise categorization of unsupervised TSAD, the fundamental consensus in current research is that the training phase does not directly use anomaly sample labels for model optimization. Unless otherwise specified, this paper adopts this overarching philosophy and collectively refers to both UTAD-I and UTAD-II as unsupervised techniques (UTAD).
In contrast, Supervised Time series Anomaly Detection (STAD) methods utilize anomaly sample labels during the training phase to optimize performance (as shown in Figure 1(c)), but such methods have received limited research attention [15, 26, 56].
Figure 1: Classification of time series anomaly detection approaches.
However, in practical applications, a small amount of anomaly sample labels can be acquired, and the purely unsupervised assumption often struggles to hold true [54]. Existing studies largely overlook the significance of utilizing these few labels in time series anomaly detection, focusing instead primarily on improvements to model architecture.
Is relentlessly pursuing complex model architecture optimization truly the best path for time series anomaly detection research? Given a limited labeling budget, how significant is the performance gain achieved by incorporating labels?
To address these critical questions, rather than pursuing marginal gains through architectural complexity, we propose a deliberately minimalist Supervised Time-series ANomaly Detection (STAND) baseline. STAND is intentionally stripped of overly complex mechanisms to serve as a pure conceptual probe, explicitly isolating and quantifying the true impact of supervisory signals. Furthermore, we establish a comprehensive unified benchmark bridging the gap between supervised and unsupervised TSAD evaluation. Through extensive empirical and theoretical analysis, we derive the following key findings and contributions:
  1. 1.
    Paradigm Shift via Empirical Evidence: We demonstrate that under a limited labeling budget, the simple STAD method (STAND) significantly outperforms existing highly complex unsupervised TSAD architectures. As illustrated in Figure 2, the STAD category consistently yields superior performance, proving our core thesis that ”Labels Matter More Than Models.”
  2. 2.
    Quantifying the Supervisory Gain: We theoretically and empirically validate that the performance improvement gained from minimal supervision (e.g., utilizing as little as 10% of the data) far outweighs the incremental gains typically achieved through unsupervised model architecture improvements.
  3. 3.
    Readability & Practicality in Real-world Scenarios: We reveal that existing UTAD methods often fail to clearly delineate anomaly segments and lack prediction consistency. In contrast, STAD methods successfully anchor anomalies while maintaining robust structural consistency, demonstrating significantly stronger practical applicability even under label noise.
  4. 4.
    A Unified Evaluation Benchmark and Library: Moving beyond mere model evaluation, we present a robust benchmark framework and an open-source library  [^1] https://github.com/EmorZz1G/STAND that successfully incorporates supervised methods into the TSAD evaluation landscape alongside unsupervised baselines, fostering future data-centric TSAD research.
Figure 2: Multi-metric performance comparison across different categories of time series anomaly detection methods.
II Related Work
II-A Algorithm-Centric Time Series Anomaly Detection
Early unsupervised Time Series Anomaly Detection (UTAD) methods were primarily based on classical machine learning techniques, such as Principal Component Analysis (PCA) [31], Local Outlier Factor (LOF) [5], and Isolation Forest (IForest) [24]. These methods transform time series data into static feature vectors using techniques like sliding window segmentation and statistical feature extraction, which are then fed into traditional anomaly detection algorithms. However, these approaches heavily depend on the quality of hand-crafted features, and their model structures are often insufficiently adapted to the dynamic nature of time series data.
With the advancement of deep learning technology, its powerful capabilities for automatic feature extraction and complex pattern modeling have been widely applied in the field of UTAD. Existing algorithm-centric research in this area can be broadly classified into two main technical paths:
Path I: Focusing on Model Architecture Improvement to Enhance Time Series Feature Extraction Capabilities. This includes methods based on improvements to classical neural networks [9, 1, 16]; Transformer-based approaches [44], which leverage attention mechanisms to improve long-range dependency modeling; Mamba-based methods [21, 39, 3], which enhance the efficiency of processing long time sequences through state space models; and Large Language Model (LLM)-based methods [55, 23], which utilize the general knowledge learned by LLMs for anomaly recognition.
Path II: Focusing on Establishing Anomaly Priors to Optimize Anomaly Determination. Examples include networks based on one-class classification and Random Forest principles [29, 42], which introduce the concept of data centricity to identify anomalies; methods based on dissimilarity learning [52, 53, 51, 45], which employ proxy strategies to learn the similarity distribution between normal and anomalous instances; and methods based on time-frequency information learning [47, 25, 18], which transform time series data into the frequency domain to capture multi-granularity anomalies.
All the aforementioned paths originate from the algorithmic level, attempting to overcome the label scarcity problem by constructing powerful models, but they fail to address the core contradiction directly from the data perspective.
II-B Data-Centric Time Series Anomaly Detection
Unlike the goal of algorithm-centric research, data-centric time series anomaly detection research revolves around optimizing the data layer to compensate for the deficiency of labels in unsupervised settings. However, due to the low prevalence of anomaly events and the high cost of labeling in most practical scenarios, studies that directly leverage high-quality anomaly labels to optimize models remain scarce [15, 19, 4]. Based on this constraint, past data-centric research on time series anomaly detection has mainly explored two optimization paths:
Path I: Pseudo-Sample-Based Data Augmentation Strategies to Expand Data Value with Limited Information. The core of this path is to substitute real labels with data transformations, creating pseudo-anomaly samples or task-oriented samples to provide learning signals for the model. This is implemented in two ways: First, unsupervised pseudo-sample generation, where normal samples are modified into anomaly samples in a label-free setting by simulating anomaly mechanisms (e.g., anomaly injection, segment shuffling, etc.). This guides the model to explore anomaly patterns and enhance detection capability [43, 38, 14]. Second, construction of task-guided samples, where, for instance, noisy time series data are treated as boundary samples between normal and anomalous states, or similar samples are generated via operations like Mixup, interpolation, or inversion to assist with model initialization [50, 51].
Path II: New Sample Generation Strategies Based on Distribution Modeling, Guiding Pattern Learning through Data Generation. This path utilizes generative models (e.g., GANs, VAEs, Diffusion Models) to learn the distribution characteristics of normal time series data, thereby generating new samples that conform to realistic patterns and reinforcing the model’s perception of normal modes [7, 20]. The generated samples can directly augment the training set, mitigating data scarcity. Simultaneously, an anomaly determination baseline is established through the detection of deviations from the learned normal distribution, forming a closed loop between data generation and anomaly detection.
Despite the fact that these studies have explored the improvement of model performance via data quality enhancement, their core focus remains restricted to the unsupervised framework. They essentially bypass the reliance on true labels through pseudo-samples or generated samples, failing to fully exploit the few anomaly labels that are practically available. This represents a shared limitation of both the algorithm-centric and current data-centric approaches: both inherently assume a completely label-free prerequisite, yet overlook the potential gain in performance from the supervisory signal provided by limited labeling.
III Methodology
To clearly contrast the core differences between various paradigms of TSAD and to highlight the value of the supervisory signal, this section first establishes the mathematical definitions for unsupervised, supervised, and our proposed STAND methods. Following this, the specific implementation of the STAND baseline is detailed.
III-A Definitions
Definition 1 (Time Series Anomaly Detection).
The multivariate time series sample is represented as $\mathbf{X}=[x_{1},x_{2},\dots,x_{T}]^{\top}\in\mathbb{R}^{T\times C}$, where $T$ is the time series length, $C$ denotes the number of variables, and $x_{t}\in\mathbb{R}^{C}$ is the multivariate observation at time $t$. The corresponding time-step labels are $\mathbf{y}=[y_{1},y_{2},\dots,y_{T}]^{\top}$, where $y_{t}=0$ indicates a normal time step and $y_{t}=1$ indicates an anomalous time step. The objective of anomaly detection is to learn a mapping function $f:\mathbb{R}^{T\times C}\rightarrow\mathbb{R}^{T}$, which outputs the anomaly score sequence for each time step, $\mathbf{S}=f(\mathbf{X})$. A preset threshold $\tau$ is then used for class determination: if $s_{t}>\tau$, time step $t$ is classified as anomalous; otherwise, it is normal.
Definition 2 (Unsupervised Time-series Anomaly Detection (UTAD)).
The core assumption of UTAD is that the proportion of anomaly samples in the training data is extremely low or the training data is entirely anomaly-free. The training set is defined as $\mathcal{D}_{\text{trn}}=\{x_{i}\}_{i=1}^{N_{\text{trn}}}$. The method’s goal is to capture the distribution characteristics of normal time series via unsupervised learning, generating training set anomaly scores $\mathbf{S}_{\text{trn}}=f(\mathbf{X}_{\text{trn}})$, and then inferring the scores $\mathbf{S}_{\text{test}}=f(\mathbf{X}_{\text{test}})$ on the test set $\mathcal{D}_{\text{test}}=\{x_{j}\}_{j=1}^{N_{\text{test}}}$, which contains anomalies.
Definition 3 (Supervised Time-series Anomaly Detection (STAD)).
STAD is applicable in scenarios where a small number of anomaly labels are available. Its training set is defined as the labeled data $\mathcal{D}_{\text{trn}}=\{(x_{i},y_{i})\}_{i=1}^{N_{\text{trn}}}$. The method’s goal is to leverage supervisory information to guide the model in learning the mapping relationship, generating training set anomaly scores $\mathbf{S}_{\text{trn}}=f(\mathbf{X}_{\text{trn}},\mathbf{y})$, and subsequently detecting anomalies in the test set $\mathcal{D}_{\text{test}}$.
III-B Proposed Method
In this section, we introduce our proposed Supervised Time-series ANomaly Detection (STAND) framework. Our core hypothesis is that even a rudimentary model, when guided by a small amount of supervision, can decisively outperform complex unsupervised architectures.
Consequently, STAND is intentionally designed with a minimalist and structurally straightforward architecture. By deliberately eschewing highly parameterized or overly intricate components (such as deep self-attention mechanisms or complex generative bottlenecks), we ensure that STAND serves as a pure proof-of-concept baseline. This minimalist design guarantees that the observed performance superiority stems entirely from the utilization of supervisory signals rather than architectural over-engineering, thereby empirically validating the significant impact of labels. The framework treats time-series anomaly detection as a sequence-to-sequence binary classification task at the timestamp level. As illustrated in Figure 3, STAND is comprised of three basic components: a Feature Embedding, a Temporal Encoding Module, and an Anomaly Scoring Module.
Figure 3: The overall architecture of the STAND  framework. An input time series is first projected into latent space by the feature embedding. The resulting latent sequence is then fed into the temporal encoding module, which captures bidirectional temporal context. Finally, an anomaly scoring module outputs anomaly scores for each time step.
III-B1 Model Architecture
Feature Embedding. The raw input at each timestamp, $x_{t}\in\mathbb{R}^{C}$, often resides in a feature space that may not be optimal for separating normal and anomalous patterns. To address this, we first employ a Feature Embedding to project the input features into a more discriminative latent space. This module enhances the model’s capacity to learn complex, non-linear relationships between the input variables before they are processed sequentially.
As defined in our implementation, this module is a simple Multi-Layer Perceptron (MLP) with two linear layers, each followed by a Gaussian Error Linear Unit (GELU) activation function [11] and Layer Normalization [2]. The process for a single timestamp $x_{t}$ can be formulated as:
$h_{t}^{(0)}=x_{t}$ (1)
$h_{t}^{(l)}=\text{LayerNorm}(\text{GELU}(\mathbf{W}^{(l)}h_{t}^{(l-1)}+b^{(l)}))$ (2)
where $l$ is the layer index in the MLP. The final output of this module is a sequence of embeddings $\mathbf{H}^{e}=[h_{1}^{e},h_{2}^{e},\dots,h_{T}^{e}]^{\top}\in\mathbb{R}^{T\times d_{\text{model}}}$, where $d_{\text{model}}$ is the hidden dimension of the model. This embedding process is applied independently to each timestamp. An option to bypass this module is included for ablation studies, in which case the input is fed directly to the next module.
Temporal Encoding Module. To capture the temporal dependencies inherent in time-series data, the sequence of embeddings $\mathbf{H}^{e}$ is passed to a Temporal Encoding Module. The choice of encoder is critical for modeling the contextual information that distinguishes an anomaly from normal behavior. We utilize a Long Short-Term Memory (LSTM) network [12], a proven and effective recurrent neural network (RNN) architecture for sequential data. LSTMs use a gating mechanism (input, forget, and output gates) to regulate information flow, enabling them to capture long-range dependencies while mitigating the vanishing gradient problem.
To provide the model with a comprehensive contextual understanding for each timestamp, we employ a bidirectional LSTM. A bidirectional LSTM processes the input sequence in both forward (from $t=1$ to $T$) and backward (from $t=T$ to $1$) directions. The hidden state for each timestamp $t$ is the concatenation of the forward hidden state $\overrightarrow{h_{t}}$ and the backward hidden state $\overleftarrow{h_{t}}$:
$\overrightarrow{h_{t}}=\text{LSTM}_{\text{fwd}}(h_{t}^{e},\overrightarrow{h_{t-1}})$ (3)
$\overleftarrow{h_{t}}=\text{LSTM}_{\text{bwd}}(h_{t}^{e},\overleftarrow{h_{t+1}})$ (4)
$h_{t}^{\text{enc}}=[\overrightarrow{h_{t}};\overleftarrow{h_{t}}]$ (5)
where $\cdot$$\cdot$] denotes concatenation. The resulting hidden state $h_{t}^{\text{enc}}\in\mathbb{R}^{2\times d_{\text{model}}}$ encapsulates information from both past and future contexts. This is particularly valuable for anomaly detection, as the nature of an event at time $t$ can often be better understood by observing subsequent data points. The module can also be configured with multiple LSTM layers to learn hierarchical temporal features. The output is a sequence of contextualized representations $\mathbf{H}^{\text{enc}}=[h_{1}^{\text{enc}},\dots,h_{T}^{\text{enc}}]^{\top}$.
Anomaly Scoring Module. Finally, the Anomaly Scoring Module maps the contextualized representation $h_{t}^{\text{enc}}$ of each timestamp to a single scalar value. This scalar represents the raw, unnormalized score (logit) indicating the likelihood of that timestamp being anomalous. This is achieved with a simple linear layer (the classifier) applied pointwise to each element of the sequence $\mathbf{H}^{\text{enc}}$:
$s_{t}=\mathbf{W}_{c}h_{t}^{\text{enc}}+b_{c}$ (6)
where $\mathbf{W}_{c}\in\mathbb{R}^{1\times(2\cdot d_{\text{model}})}$ and $b_{c}\in\mathbb{R}$ are the weight and bias of the linear classifier. The model outputs a sequence of anomaly logits $\mathbf{S}=[s_{1},s_{2},\dots,s_{T}]^{\top}\in\mathbb{R}^{T}$. These logits can be converted into probabilities using the sigmoid function, $p_{t}=\sigma(s_{t})$, for interpretation or thresholding.
III-B2 Training Objective
Given the availability of timestamp-level labels $\mathbf{y}=[y_{1},\dots,y_{T}]^{\top}$, we can train the STAND model in an end-to-end supervised fashion. We frame the task as a binary classification problem for each timestamp. The model is optimized by minimizing the Binary Cross-Entropy (BCE) loss between the predicted scores and the ground-truth labels. To improve numerical stability, we introduce a sigmoid layer into the computation of BCE loss. The loss for a single time-series sample $(\mathbf{X},\mathbf{y})$ is calculated as:
$\mathcal{L}(\mathbf{S},\mathbf{y})=-\frac{1}{T}\sum_{t=1}^{T}[y_{t}\cdot\log(\sigma(s_{t}))+(1-y_{t})\cdot\log(1-\sigma(s_{t}))]$ (7)
where $s_{t}$ is the output logit from the model for timestamp $t$, $y_{t}\in\{0,1\}$ is the corresponding ground-truth label, and $\sigma(\cdot)$ is the sigmoid function. The total loss is averaged over all samples in a mini-batch. The model parameters are updated via backpropagation using a standard gradient-based optimizer such as Adam [17]. The complete training and inference procedures are summarized in Algorithm 1.
Algorithm 1 Training and Inference with STAND
1:Input: Training set $\mathcal{D}_{\text{trn}}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N_{\text{trn}}}$, test sequence $\mathbf{X}_{\text{test}}$, learning rate $\eta$, number of epochs $E$.
2:Output: Anomaly scores $\mathbf{S}_{\text{test}}$ for $\mathbf{X}_{\text{test}}$.
3:
4:procedure Train($\mathcal{D}_{\text{trn}},\eta,E$)
5:  Initialize STAND model parameters $\theta$.
6:  for epoch = 1 to $E$ do
7:   for each batch $(\mathbf{X},\mathbf{y})$ in $\mathcal{D}_{\text{trn}}$ do
8:     # Forward pass
9:     Compute anomaly logits: $\mathbf{S}\leftarrow\text{STAND}(\mathbf{X};\theta)$.
10:     # Compute loss
11:     Compute loss $\mathcal{L}$.
12:     # Backward pass and optimization
13:     Update parameters: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$.
14:   end for
15:  end for
16:  return Trained parameters $\theta$.
17:end procedure
18:
19:procedure Inference($\mathbf{X}_{\text{test}},\theta$)
20:  Load STAND model with trained parameters $\theta$.
21:  Compute anomaly logits: $\mathbf{S}_{\text{test}}\leftarrow\text{STAND}(\mathbf{X}_{\text{test}};\theta)$.
22:  return Anomaly scores $\mathbf{S}_{\text{test}}$.
23:end procedure
IV Theoretical Analysis
To theoretically substantiate the proposed STAND framework and our core motivation, we rigorously examine its scalability, optimization stability, and sample complexity.
IV-A Scalability Analysis
The highly scalable nature of STAND is rooted in its computational efficiency, a direct consequence of its streamlined architecture, which eschews the quadratic complexity often associated with canonical attention mechanisms.
Let the input time series be denoted by $\mathbf{X}\in\mathbb{R}^{T\times C}$, where $T$ represents the temporal length and $C$ is the number of variables. We denote the hidden dimension of the model as $d$ and the number of layers in the Temporal Encoding Module as $L$.
The complexity of the Feature Embedding module, which projects inputs via a multi-layer perceptron, is computed pointwise. For the entire sequence, the computational cost $\mathcal{C}_{\text{embed}}$ is formulated as:
$\mathcal{C}_{\text{embed}}=\mathcal{O}(T\cdot C\cdot d)$ (8)
Following the embedding, the bidirectional LSTM employed in the Temporal Encoding Module processes the sequence recurrently. The state update for a single LSTM cell involves matrix-vector multiplications proportional to the square of the hidden dimension. Consequently, for a sequence of length $T$ and a bidirectional configuration across $L$ layers, the temporal processing cost $\mathcal{C}_{\text{temp}}$ is bounded by:
$\mathcal{C}_{\text{temp}}=\mathcal{O}(T\cdot d^{2}\cdot L)$ (9)
The Anomaly Scoring Module acts as a linear projection from the hidden space to the scalar anomaly score, contributing a cost $\mathcal{C}_{\text{score}}$:
$\mathcal{C}_{\text{score}}=\mathcal{O}(T\cdot d)$ (10)
By aggregating these components, the total time complexity per training epoch, denoted as $\mathcal{T}_{\text{total}}$, is derived as:
$\mathcal{T}_{\text{total}}=\mathcal{C}_{\text{embed}}+\mathcal{C}_{\text{temp}}+\mathcal{C}_{\text{score}}\approx\mathcal{O}(T(Cd+Ld^{2}))$ (11)
This formulation highlights a critical theoretical advantage regarding scalability for big data applications: the computational complexity scales strictly linearly with the sequence length $T$, making STAND highly efficient for processing massive, long-term time series streams. In stark contrast, state-of-the-art unsupervised methods relying on self-attention mechanisms, such as Anomaly Transformer, exhibit a quadratic complexity $\mathcal{T}_{\text{Trans}}$:
$\mathcal{T}_{\text{Trans}}\approx\mathcal{O}(T^{2}\cdot d)$ (12)
This makes $\mathcal{T}_{\text{Trans}}\gg\mathcal{T}_{\text{total}}$ for long sequence inputs ($T\gg d$).
Regarding memory constraints, the space complexity consists of parameter storage $\mathcal{M}_{\text{param}}$ and runtime activation memory $\mathcal{M}_{\text{run}}$. The parameter space is dominated by the LSTM weights:
$\mathcal{M}_{\text{param}}=\mathcal{O}(L\cdot d^{2}+C\cdot d)$ (13)
Meanwhile, the runtime memory required for storing hidden states for backpropagation scales strictly linearly:
$\mathcal{M}_{\text{run}}=\mathcal{O}(T\cdot d\cdot L)$ (14)
Similar to the computational cost, this linear memory dependence offers a crucial scalability advantage over Transformer-based models, which typically require $\mathcal{O}(T^{2})$ memory for attention matrices. This ensures that STAND maintains a manageable memory footprint even when processing massive data streams, further cementing its suitability for large-scale industrial deployments.
IV-B Convergence Analysis
Beyond efficiency, the utilization of supervision signals introduces a stable gradient flow that facilitates convergence. We formulate the learning objective as an Empirical Risk Minimization (ERM) problem.
Let $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ be the training dataset. The objective function $J(\theta)$ is defined as the average Binary Cross-Entropy loss over the dataset:
$\small J(\theta)=-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left[y_{i,t}\log(\sigma(s_{i,t}))+(1-y_{i,t})\log(1-\sigma(s_{i,t}))\right]$ (15)
where $\theta$ represents the set of learnable parameters, and $s_{i,t}$ is the logit output at time $t$.
To analyze convergence, we assume that the gradient of the objective function, $\nabla J(\theta)$, is $\mathcal{K}$-Lipschitz continuous. This implies that for any two parameter vectors $\theta_{1},\theta_{2}$, the following inequality holds:
$\|\nabla J(\theta_{1})-\nabla J(\theta_{2})\|\leq\mathcal{K}\|\theta_{1}-\theta_{2}\|$ (16)
where $\mathcal{K}>0$ is the Lipschitz constant.
Consider the parameter update rule using standard Gradient Descent with a learning rate $\eta$:
$\theta_{k+1}=\theta_{k}-\eta\nabla J(\theta_{k})$ (17)
According to the Descent Lemma for Lipschitz continuous functions, the objective value at the next iteration is bounded by:
$J(\theta_{k+1})\leq J(\theta_{k})+\langle\nabla J(\theta_{k}),\theta_{k+1}-\theta_{k}\rangle+\frac{\mathcal{K}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$ (18)
Substituting the update rule into this inequality yields:
$J(\theta_{k+1})$ $\leq J(\theta_{k})-\eta\|\nabla J(\theta_{k})\|^{2}+\frac{\mathcal{K}\eta^{2}}{2}\|\nabla J(\theta_{k})\|^{2}$ (19)
$=J(\theta_{k})-\eta\left(1-\frac{\mathcal{K}\eta}{2}\right)\|\nabla J(\theta_{k})\|^{2}$
Assuming the learning rate is chosen such that $\eta<\frac{2}{\mathcal{K}}$, the term $\eta(1-\frac{\mathcal{K}\eta}{2})$ is strictly positive. This guarantees that the sequence of loss values $\{J(\theta_{k})\}$ is non-increasing.
By summing the inequality over $K$ iterations and rearranging terms, we obtain:
$\sum_{k=0}^{K}\eta\left(1-\frac{\mathcal{K}\eta}{2}\right)\|\nabla J(\theta_{k})\|^{2}\leq J(\theta_{0})-J(\theta_{K+1})$ (20)
Since the Binary Cross-Entropy loss is bounded below by 0, i.e., $J(\theta)\geq 0$, the right-hand side is bounded by $J(\theta_{0})$. Taking the limit as $K\to\infty$, the series converges:
$\lim_{K\to\infty}\sum_{k=0}^{K}\|\nabla J(\theta_{k})\|^{2}<\infty$ (21)
This implies that the gradient norm must approach zero:
$\lim_{k\to\infty}\|\nabla J(\theta_{k})\|=0$ (22)
While this represents a standard descent argument and global optimality is not theoretically guaranteed due to the non-convex nature of the recurrent architecture, Eq. (14) rigorously confirms that STAND converges stably to a critical point. We include this analysis to draw a deliberate contrast: unlike state-of-the-art UTAD methods relying on adversarial training (e.g., GANs) or complex reconstruction bottlenecks—which are notoriously prone to mode collapse and training instability—our supervised objective maintains standard, reliable convergence properties. This theoretical stability translates directly into the empirical robustness and prediction consistency observed in our evaluations.
IV-C Sample Complexity and the Benefit of Supervision
To rigorously substantiate our empirical finding that labels matter more than models, we analyze the intrinsic learning difficulty of Unsupervised Time-series Anomaly Detection (UTAD) versus Supervised Time-series Anomaly Detection (STAD) from the perspective of Statistical Learning Theory [37, 30]. Specifically, we formulate their respective risk functions and compare the required sample complexities to achieve a bounded generalization error $\epsilon$.
Unsupervised Complexity (UTAD): UTAD methods fundamentally rely on modeling the intrinsic distribution of normal data within a high-dimensional feature space $\mathbb{R}^{C}$. Let $p^{*}(x)$ be the true probability density function of the normal data, and $\hat{p}(x)$ be the estimated density function. In this paradigm, identifying anomalies equates to non-parametric density estimation, where the learning objective is to minimize the $L_{\infty}$ risk:
$\mathcal{R}_{\text{UTAD}}(\hat{p})=\sup_{x\in\mathbb{R}^{C}}|\hat{p}(x)-p^{*}(x)|$ (23)
According to minimax lower bounds in non-parametric statistics [34, 32], estimating a density function that is $s$-times differentiable in $\mathbb{R}^{C}$ with an expected error bound $\mathbb{E}[\mathcal{R}_{\text{UTAD}}(\hat{p})]\leq\epsilon$ requires a sample size $N_{\text{UTAD}}$ satisfying:
$N_{\text{UTAD}}=\Omega\left(\epsilon^{-\frac{C}{s}}\right)$ (24)
This exponential dependence on the variable dimension $C$ manifests the curse of dimensionality. Consequently, as the number of time-series channels increases (e.g., $C=122$ in the WADI dataset), purely unsupervised methods demand an exponentially larger normal dataset to accurately map the data manifold. This theoretical bottleneck explains why current algorithm-centric UTAD methods, despite employing complex architectures, are highly susceptible to overfitting background noise and producing false positives.
Supervised Complexity (STAD): Conversely, STAD reframes the objective from high-dimensional density estimation to a discriminative binary classification task. Let $\mathcal{H}$ be the hypothesis class of the supervised model (e.g., the neural architecture of STAND) with a Vapnik-Chervonenkis (VC) dimension of $d_{\text{VC}}(\mathcal{H})$. For a specific classifier $h\in\mathcal{H}$ and label sequence $\mathbf{y}$, the expected risk $R(h)$ and empirical risk $\hat{R}(h)$ are defined as:
$R(h)=\mathbb{E}_{(x,y)\sim\mathcal{P}_{X,Y}}[\ell(h(x),y)]$ (25)
$\hat{R}(h)=\frac{1}{N_{\text{STAD}}}\sum_{i=1}^{N_{\text{STAD}}}\ell(h(x_{i}),y_{i})$ (26)
By leveraging the explicit labels, the algorithm learns a decision boundary via Empirical Risk Minimization. Under the agnostic Probably Approximately Correct (PAC) learning framework [36], the generalization gap is bounded with a probability of at least $1-\delta$ by:
$\sup_{h\in\mathcal{H}}|R(h)-\hat{R}(h)|\leq\mathcal{O}\left(\sqrt{\frac{d_{\text{VC}}(\mathcal{H})+\ln(1/\delta)}{N_{\text{STAD}}}}\right)$ (27)
To ensure the generalization error is strictly bounded by $\epsilon$, the required labeled sample size $N_{\text{STAD}}$ converges to:
$N_{\text{STAD}}=\mathcal{O}\left(\frac{d_{\text{VC}}(\mathcal{H})+\ln(1/\delta)}{\epsilon^{2}}\right)$ (28)
This bound reveals that the sample complexity of STAD grows only polynomially with respect to $1/\epsilon$ and linearly with the model’s capacity $d_{\text{VC}}(\mathcal{H})$, entirely circumventing the exponential penalty $\epsilon^{-C/s}$.
Remark: The theoretical divergence between Eq. (24) and Eq. (28) mathematically formalizes our core motivation: introducing supervisory signals collapses an exponentially hard distribution-matching problem into a polynomially solvable discrimination problem. This rigorously explains the phenomena observed in our experiments (Sec. V-D), where a simple supervised baseline trained on a drastically reduced data budget (e.g., 20% of the labeled data) can decisively outperform state-of-the-art unsupervised architectures attempting to learn the full input distribution.
V Experiments
V-A Task Designs
To validate the motivation of this paper, we establish the following experimental tasks:
  1. 1.
    Comparison of Unsupervised and Supervised Methods: To analyze the performance of the two method categories, we follow the conventional approach and test unsupervised methods under the unsupervised condition. Constrained by the scarcity of labels, we use half of the labeled data as the training set for supervised methods, and test them on the original test set.
  2. 2.
    Analysis of Supervisory Gain: We further reduce the amount of labeled time series used for training (utilizing as little as 10% of the data) and test on the original test set. We use three subsets of the PSM dataset (SP1-SP3) with increasing anomaly counts (from least to most) for this analysis. The specific dataset partitioning method and data characteristics are introduced in Sec. V-B3.
  3. 3.
    Robustness Analysis to Label Noise: To evaluate the reliability of STAD methods under imperfect annotation conditions, we simulate realistic annotator errors by injecting varying proportions of point and segment noise into the training labels, thereby assessing the models’ resilience to corrupted supervisory signals.
  4. 4.
    Effectiveness Analysis of STAND: To verify the effectiveness of our proposed STAND method, we conduct ablation experiments on its key components.
  5. 5.
    Sensitivity Analysis of STAND: To explore the influence of parameters on STAND, we perform an analysis of important hyperparameters.
  6. 6.
    Visualization Analysis: To further analyze the advantages of supervised methods, we conduct a visualization analysis of the outputs from different models.
V-B Task Setups
V-B1 Metrics
To evaluate the global anomaly detection capability of the different models, we adopt the point-wise evaluation metrics, F1 score and Area Under the Receiver Operating Characteristic Curve (AUC-ROC).
To assess the models’ event-level anomaly detection capability, we introduce the Aff-F1, UAff-F1 [51], and Volume Under the Surface for the Precision-Recall surface (VUS-PR) [27] scores. Aff-F1 evaluates the interval-level F1-score by utilizing interval membership to quantify the distance between predicted results and ground truth labels. Furthermore, UAff-F1 is an unbiased refinement of Aff-F1, placing greater emphasis on the precision of the model’s detection. VUS-PR is the area under the event-level Precision-Recall curve, which effectively assesses a model’s ability to detect anomalous events.
Furthermore, we introduce Confidence-Consistency Evaluation (CCE) to quantify the prediction consistency of the models [49]. This metric not only considers the accuracy of the anomaly scores but also accounts for local and global prediction uncertainty. It has been theoretically proven to be more robust and is particularly suitable for scenarios where anomaly labels or anomaly scores are noisy.
V-B2 Baselines
The UTAD-I methods consist of traditional unsupervised anomaly detection approaches, covering Random Guess (Random), a method that predicts each time point as anomalous with a 50% probability and serves as an intuitive baseline to contrast the performance gaps between other methods and a random strategy, Isolation Forest (IForest), Local Outlier Factor (LOF), Principal Component Analysis (PCA), Histogram-based Outlier Score (HBOS), K-Nearest Neighbors (KNN), and K-Means Clustering (KMeans). These methods rely on classical statistical characteristics, distance metrics, or clustering mechanisms to identify anomalies without requiring labeled training data.
The UTAD-II methods encompass a broader range of techniques, including traditional kernel methods, deep learning models, and Transformer-based approaches, specifically including One-Class Support Vector Machine (OCSVM), Autoencoder (AE), Convolutional Neural Network (CNN), Long Short-Term Memory Network (LSTM), TranAD [35], USAD [1], Omni [33], Anomaly Transformer (A.T.) [44], TimesNet [40], M2N2 [16], LFTSAD [6], and CATCH [41]. This category integrates shallow kernel-based learning, deep neural networks (for spatial-temporal feature extraction), and state-of-the-art Transformer architectures (for capturing long-range dependencies in time series).
The STAD methods are composed of traditional supervised learning classification models, including Random Forest (RF), Support Vector Machine (SVM), AdaBoost, Extra-Trees Classifier (ExtraTrees), and Histogram-based Gradient Boosting Classification Tree (LightGBM). These models are trained on labeled normal/anomalous samples to learn discriminative patterns for direct anomaly classification.
Implementation Details:
  • All methods in UTAD-I, as well as OCSVM, AE, CNN, LSTM, TranAD, USAD, Omni, Anomaly Transformer (A.T.), and TimesNet in UTAD-II, adopt the implementations provided by the TSB-AD benchmark library [^2] https://github.com/TheDatumOrg/TSB-AD. TSB-AD is a widely recognized time-series anomaly detection benchmark that ensures consistent experimental settings (e.g., hyperparameter ranges, data preprocessing) across different models, enhancing the reproducibility of our comparisons.
  • For M2N2, LFTSAD, and CATCH in UTAD-II, we implement the models based on the official open-source code released by the authors of their original papers. We adhere to the recommended hyperparameter configurations, network architectures, and training protocols reported in the literature to maintain the authenticity of the baseline performances.
  • All methods in STAD are implemented based on the Scikit-learn library [^3] https://scikit-learn.org/, a mature and extensively validated machine learning toolkit in the Python ecosystem. The Scikit-learn implementations are chosen for their stability, efficiency, and widespread adoption in academic and industrial research, ensuring reliable baseline results.
V-B3 Datasets
To evaluate the performance and scalability of different algorithms in real-world large-scale scenarios, we utilize five well-known multivariate time series anomaly detection datasets collected from complex industrial systems and sensor networks: PSM, SWaT, WADI, Swan, and Water [45, 51]. Notably, datasets like WADI encompass up to 122 highly correlated channels and hundreds of thousands of timestamps, representing typical industrial big data environments. The characteristics of these datasets are summarized in Table I.
For Task 2, to further investigate the gain conferred by supervision, we select PSM as the base dataset. We partition the original test set’s initial portion to serve as the training set. To ensure a fair evaluation and prevent train-test contamination, the evaluation is performed exclusively on the Remaining portion of the data (as denoted in Table II). This process results in the three disjoint PSM subsets shown in Table II (i.e., SP1, SP2, and SP3). It is important to note that only PSM satisfies the requirement that anomalies are present throughout the entire test set. The anomaly distribution patterns in the other datasets lack sufficient regularity, making it difficult to partition the dataset—specifically, splitting the entire test set into two continuous segments. Hence, we exclusively chose PSM as the base dataset for this task.
Specifically, we perform the dataset partitioning operation as follows: We first select a specific Split Threshold, such as $10\%$ or $20\%$, which indicates that the anomaly label ratio in the training set must be at least greater than this threshold. For instance, SP1 selects $10\%$ as the split threshold. Its resulting training set has an anomaly ratio of $13.35\%$ and contains $79,057$ remaining time steps of data.
TABLE I: Characteristics of the five real-world datasets. #Channel denotes the number of variables in the time series data. #Train and #Test represent the number of time steps in the training and test sets, respectively.
Dataset #Channel #Train #Test Anomaly Rate
PSM 25 132481 87841 27.76%
SWaT 50 495000 449919 12.14%
WADI 122 241921 34561 5.74%
Swan 38 60000 60000 32.60%
Water 9 69260 69261 1.05%
TABLE II: Characteristics of the PSM subsets.
Dataset Split Threshold #Train Training Anomaly Rate #Remaining Remaining Anomaly Rate #Overall Overall Anomaly Rate
SP1 10.00% 8784 13.35% 79057 29.36% 87841 27.76%
SP2 20.00% 20182 20.00% 67659 30.07%
SP3 30.00% 35136 30.73% 52705 25.77%
V-C Comparison of Unsupervised and Supervised Methods
Table III presents the results for a selection of the better-performing UTAD-I, UTAD-II, and STAD methods; the full results are available in Appendix A. It is evident that unsupervised methods (UTAD-I and UTAD-II) are rarely the top-performing approaches across all datasets. Furthermore, they often exhibit suboptimal overall performance on specific metric categories, which suggests limitations in their practical utility. For instance, recent methods like M2N2 and CATCH show low CCE scores across all five datasets, while methods such as Omni and A.T. achieve low Aff-F1 and UAff-F1 scores on PSM, SWaT, and Swan.
Figure 2 illustrates the average performance of different method categories across multiple datasets and metrics. Although UTAD-II, which incorporates the “normal-only” prior, achieves slightly better average results than UTAD-I, their average performance is not even competitive with simple baselines like PCA.
In contrast to UTAD-I and UTAD-II, STAD methods show a substantial improvement in overall integrated performance, elevating the score by $28.83$ and $24.02$ respectively. Among all methods, our proposed STAND baseline demonstrates the best comprehensive performance. This strongly suggests that STAD methods possess significantly greater potential compared to UTAD approaches. Additionally, the average CCE score for all UTAD models does not exceed $10$, which is far lower than that of the STAD category. This indicates poor prediction consistency for UTAD methods, leading to insufficient practical availability in real-world scenarios.
This comparison experiment was initially designed to maintain consistency with previous related work. However, conducting it under supervised conditions might introduce an element of unfair comparison. Therefore, we further adopt a fairer setting (i.e., evaluating models only on the entirely unseen test set) to compare UTAD and STAD methods in Appendix B. The analysis shows that regardless of the experimental setting, the overall performance of STAD is superior to both UTAD-I and UTAD-II, and the performance of STAND significantly surpasses the current state-of-the-art methods.
Finally, STAND exhibits relatively stronger performance on SWaT and WADI (datasets with larger data volume) compared to other STAD models, but performs less effectively on smaller datasets. This observation suggests that when the data volume is small but a limited number of labels are available, classical STAD methods are a better choice; however, when the data volume is large, the deep learning-based STAND can achieve superior results.
TABLE III: Performance comparison of the three method categories across five datasets and six metrics. Bold indicates the best performance, and underlining indicates the second-best performance. Methods highlighted in gray are UTAD-I, methods highlighted in green are UTAD-II, and methods highlighted in orange represent STAD methods.
Dataset PSM SWaT WADI
Metric CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR
Random -0.83 8.51 67.99 5.53 50.10 32.99 -0.48 7.07 68.96 -3.13 49.96 12.77 -0.07 4.90 69.16 -6.51 49.80 8.47
IForest -5.40 3.75 49.60 -1.69 46.76 29.96 -2.02 6.02 62.11 -9.86 32.74 10.37 7.83 10.99 61.99 -11.06 74.70 17.17
PCA 6.56 18.77 57.93 52.38 74.13 55.01 4.52 48.84 66.17 30.80 89.73 61.73 14.39 39.07 81.74 58.92 81.07 37.04
HBOS -3.44 8.52 44.79 -18.88 49.90 33.23 9.65 10.38 71.49 24.54 75.08 26.34 9.44 9.86 67.40 6.38 73.97 18.50
KMeans 0.12 7.85 60.05 44.55 38.64 30.14 4.79 19.49 78.12 46.24 33.06 18.18 6.46 12.72 66.56 28.87 64.70 13.79
AE 7.78 5.19 56.50 51.07 56.50 34.90 16.25 17.87 74.95 40.34 84.69 45.08 15.04 27.00 74.26 42.28 64.38 16.65
CNN 0.57 26.33 69.16 60.48 61.26 52.52 29.04 58.34 5.46 5.46 86.77 60.84 2.39 24.20 78.57 43.04 66.08 24.00
LSTM 0.60 26.86 66.48 57.44 63.88 51.85 29.54 58.32 5.41 5.41 86.63 68.63 2.47 23.98 78.70 43.07 70.80 25.67
TranAD 0.23 20.09 63.50 58.40 60.54 46.85 29.78 58.34 5.34 5.34 88.47 63.50 2.60 23.23 77.80 40.01 72.76 30.27
USAD 1.49 21.39 46.14 44.90 61.53 46.62 29.79 58.34 5.33 5.33 88.71 63.04 8.36 23.01 77.83 39.86 73.60 28.02
Omni 3.93 21.78 23.44 23.04 64.11 44.55 29.80 58.34 5.31 5.31 88.85 59.30 10.84 39.02 72.07 52.59 79.13 38.82
A.T. 1.02 18.20 48.73 42.56 59.13 43.46 29.31 58.34 5.35 5.35 74.39 47.13 3.88 28.02 74.06 22.00 61.97 20.27
M2N2 -0.07 15.26 79.11 47.11 63.35 48.17 0.01 23.99 75.65 42.15 81.15 33.29 1.69 20.75 74.89 22.97 64.38 22.25
CATCH 0.62 16.85 84.85 62.92 60.65 51.99 0.02 9.33 73.41 23.84 29.78 11.87 0.26 13.64 74.92 31.67 69.76 21.18
RF 44.33 64.86 76.49 69.43 90.77 80.86 61.42 56.73 86.06 85.08 95.92 89.24 61.83 87.71 90.38 88.91 97.58 91.77
SVM 15.66 29.80 22.34 21.81 91.93 80.46 26.95 35.33 81.49 80.25 93.29 88.62 32.21 80.41 92.49 84.41 93.89 87.95
AdaBoost 14.46 30.48 29.28 28.95 84.99 75.80 18.69 52.01 50.13 42.59 86.71 41.43 24.53 78.40 83.81 78.44 91.84 75.43
ExtraTrees 50.28 66.16 78.92 73.20 92.82 83.89 69.03 53.87 81.26 81.22 97.01 92.69 63.07 84.94 87.16 79.12 99.10 95.48
LightGBM 44.21 30.53 12.18 12.18 87.36 77.24 66.51 48.82 75.89 70.81 94.87 80.76 66.43 80.95 91.93 89.76 96.89 92.03
STAND 46.94 30.33 29.65 29.35 90.74 83.66 71.48 57.95 87.84 87.42 92.02 87.14 65.24 82.03 89.93 84.07 91.72 87.08
Dataset Swan Water Average
Metric CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR
Random -0.27 8.65 26.68 -0.69 49.79 83.90 0.72 1.53 63.75 -3.53 50.32 3.57 -0.18 6.13 59.31 -1.67 49.99 28.34
IForest 5.85 18.18 11.56 5.15 66.30 85.11 12.35 8.11 60.59 19.21 79.10 10.09 3.72 9.41 49.17 0.35 59.92 30.54
PCA 1.55 24.75 2.21 2.00 59.62 93.45 11.38 18.03 62.58 47.28 90.68 17.92 7.68 29.89 54.12 38.28 79.05 53.03
HBOS 15.92 18.01 23.08 20.14 78.82 88.77 16.45 8.16 65.15 9.51 82.93 11.87 9.60 10.99 54.38 8.34 72.14 35.74
KMeans -7.41 6.43 11.18 6.27 30.04 81.24 28.38 17.89 72.45 46.33 92.85 24.01 6.47 12.88 57.67 34.45 51.86 33.47
AE 5.95 18.91 21.89 20.80 60.01 88.57 1.76 5.68 54.44 22.81 50.71 2.36 9.36 14.93 56.41 35.46 63.26 37.51
CNN 0.56 24.84 15.92 15.78 72.99 93.71 2.62 3.96 44.78 9.19 54.04 2.55 7.04 27.53 42.78 26.79 68.23 46.72
LSTM 0.38 23.98 27.41 27.11 75.24 92.63 2.62 8.63 54.09 12.65 60.35 5.84 7.12 28.36 46.42 29.14 71.38 48.92
TranAD 0.08 15.94 18.57 5.76 65.01 91.96 0.32 6.20 50.82 13.96 52.11 6.51 6.60 24.76 43.21 24.69 67.78 47.82
USAD 0.54 20.85 11.95 11.42 61.81 92.27 3.21 9.35 41.21 13.84 71.66 6.27 8.68 26.59 36.49 23.07 71.46 47.25
Omni 1.86 24.44 0.26 -0.25 68.55 94.25 3.23 9.30 41.19 5.74 74.37 6.44 9.93 30.58 28.45 17.29 75.00 48.67
A.T. 0.02 3.45 6.64 -5.31 50.27 84.22 1.91 3.29 71.44 30.10 65.94 5.02 7.23 22.26 41.24 18.94 62.34 40.02
M2N2 0.42 26.50 0.03 0.03 77.54 96.62 0.65 22.13 84.87 63.64 91.91 37.77 0.54 21.73 62.91 35.18 75.67 47.62
CATCH 0.22 25.60 3.05 3.03 61.55 94.17 0.88 18.98 81.54 54.51 91.28 27.36 0.40 16.88 63.55 35.19 62.61 41.31
RF 62.27 45.36 81.28 80.84 96.39 97.94 70.84 39.31 81.52 78.35 93.55 55.41 50.75 37.69 71.85 66.12 92.29 75.20
SVM 21.99 25.79 51.55 51.29 89.07 97.04 51.56 32.48 98.42 96.81 98.55 82.80 21.22 28.20 61.82 53.88 86.47 67.16
AdaBoost 27.03 25.56 41.22 41.03 80.13 92.86 47.29 32.58 92.36 82.25 98.83 88.42 23.21 29.25 53.55 44.76 82.25 61.13
ExtraTrees 57.48 41.98 74.27 73.10 95.58 97.51 92.52 35.80 98.42 96.62 99.76 93.93 50.19 34.57 69.58 60.87 91.47 73.90
LightGBM 67.56 25.14 46.38 46.04 94.28 97.34 71.07 73.33 82.77 79.06 94.11 54.01 47.94 32.42 63.68 51.93 88.72 68.99
STAND 55.53 25.57 46.98 46.65 84.76 97.19 84.44 34.01 96.61 93.41 99.62 95.38 64.73 45.98 70.20 68.18 91.77 90.09
V-D Analysis of Supervisory Gain
As shown in Table IV, the incorporation of supervisory signals yields substantial performance improvements for STAD models. We can draw several key observations from these results.
First, as the availability of labeled data increases from SP1 to SP3, the performance of supervised methods demonstrates a clear and consistent upward trajectory. Specifically, STAND’s average score improves dramatically from $37.09$ on SP1 to $46.91$ on SP3. Its average rank also ascends from $2.8$ to $1.8$. This highlights the model’s exceptional capability to translate even incremental additions of supervisory information into significant performance gains.
Second, the results reveal a distinct difference among unsupervised methods based on data volume. When the training data is extremely limited (e.g., SP1), traditional UTAD-I methods like PCA actually perform remarkably well. In fact, PCA achieves a highly competitive average rank of $2.2$ on SP1. Conversely, deep learning-based UTAD-II methods, such as CATCH, require massive amounts of data to model normal distributions effectively. As a result, they struggle in low-data regimes. Even as the data volume increases in SP2 and SP3, their overall performance still falls short of the top-tier methods.
Finally, unsupervised methods like CATCH and PCA exhibit stagnant or fluctuating performance across the data subsets. This is because they cannot directly optimize using anomaly labels. In contrast, STAD methods consistently capitalize on the expanded label set. In both SP2 and SP3, STAND achieves the highest average scores ($45.95$ and $46.91$, respectively) and the best average ranks ($1.6$ and $1.8$). This compelling empirical evidence robustly validates our core hypothesis: allocating resources to obtain a small amount of anomaly labels yields far greater performance dividends than relentlessly increasing unsupervised model complexity.
TABLE IV: Performance of all supervised methods across the three PSM subsets (SP1–SP3).
Source Model CCE F1 UAff-F1 AUC-ROC VUS-PR Avg. Score Avg. Rk.
SP1 PCA 5.47 18.17 53.51 73.35 55.78 41.25 2.2
CATCH 0.59 16.89 55.80 62.88 52.47 37.73 4.4
RF 4.64 9.89 34.29 60.37 41.24 30.09 6.8
SVM 6.09 11.69 31.52 63.88 45.37 31.71 5.6
AdaBoost 6.76 15.61 31.95 71.18 49.58 35.02 3.4
ExtraTrees 6.33 16.26 30.89 66.19 47.01 33.34 4.6
LightGBM 4.71 10.15 17.20 66.20 45.81 28.81 6.2
STAND 14.83 19.68 29.74 68.17 53.01 37.09 2.8
SP2 PCA 5.51 18.71 52.88 74.29 57.11 41.70 4.2
CATCH 0.97 17.34 59.94 63.13 54.86 39.25 5.6
RF 10.27 17.00 22.77 71.47 51.47 34.60 6.2
SVM 10.21 17.19 31.41 76.14 55.27 38.04 5.2
AdaBoost 7.85 16.91 22.64 69.73 46.39 32.70 7.4
ExtraTrees 11.19 23.12 39.60 77.26 59.98 42.23 2.4
LightGBM 13.16 21.12 35.69 76.53 57.11 40.72 3.4
STAND 19.76 21.81 42.67 80.48 65.01 45.95 1.6
SP3 PCA 11.09 22.37 51.73 79.03 55.28 43.90 3.8
CATCH 2.60 15.73 53.20 60.61 45.90 35.61 6.4
RF 12.75 15.93 15.97 72.74 50.62 33.60 5.4
SVM 12.25 22.13 52.09 80.42 57.39 44.86 2.8
AdaBoost 10.77 17.95 27.34 71.87 43.73 34.33 6.8
ExtraTrees 11.55 23.60 46.20 78.33 59.07 43.75 3.2
LightGBM 11.43 16.20 35.70 72.06 47.56 36.59 5.8
STAND 24.47 26.54 42.82 80.66 60.09 46.91 1.8
V-E Robustness to Label Noise
In practical industrial environments, acquiring flawlessly accurate anomaly labels is frequently unattainable due to annotator fallibility or inherent system ambiguities. To address potential concerns regarding the reliance of STAD methods on pristine annotations, we empirically investigate the resilience of these models against label noise. We formulate two distinct noise injection paradigms on the PSM dataset to emulate realistic annotation errors:
  1. 1.
    Point Noise: We randomly sample a specified proportion $\rho$ of the labels across the entire training sequence and invert their respective classes (i.e., normal to anomalous, and vice versa). This mechanism simulates sporadic, independent annotation inaccuracies.
  2. 2.
    Segment Noise: We randomly select contiguous segments of normal time steps and reassign their labels to anomalous until the cumulative number of inverted steps reaches the target proportion $\rho$. This paradigm emulates systematic errors, wherein an annotator might misclassify an entire operational phase.
We systematically vary the noise ratio $\rho\in\{0.0,0.01,0.03,0.05,0.07,0.1\}$ and benchmark the STAD models alongside the SOTA UTAD-II method, CATCH. The comparative performance is illustrated in Figure 4.
Figure 4: Performance of supervised models under varying ratios of Point Noise (solid lines) and Segment Noise (dashed lines) on the PSM dataset. The gray dashed line represents the performance of the UTAD-II SOTA method, CATCH, which does not utilize labels and thus remains constant.
Based on the quantitative evaluations, we deduce the following critical observations:
  • Sustained Superiority Under Substantial Noise: Although the performance of all STAD models inherently degrades as the noise ratio escalates, robust methods such as ExtraTrees and STAND preserve a decisive advantage over the SOTA unsupervised baseline (CATCH) across the majority of metrics, even at a severe noise ratio of 10%. For instance, under Segment Noise at $\rho=0.1$, STAND achieves a CCE of $23.47$, an AUC of $80.91$, and a VUS-PR of $61.27$, which still substantially surpass the performance of CATCH ($0.62$, $60.65$, and $51.99$, respectively).
  • Differential Impact of Noise Paradigms: Traditional approaches like Random Forest (RF) manifest pronounced sensitivity to Point Noise, evidenced by a precipitous decline in point-wise metrics such as F1 and Aff-F1. Conversely, Segment Noise exerts a more attenuated degradative impact on these models. Notably, ExtraTrees exhibits remarkable resilience to both noise modalities, consistently maintaining top-tier global performance.
  • Robustness of STAND: The proposed STAND baseline yields stable degradation trajectories, particularly concerning global evaluation metrics (e.g., AUC) and structural consistency metrics (e.g., CCE).
    This empirical evidence suggests that learning a parametric mapping via deep sequence models affords inherent noise tolerance, thereby mitigating severe overfitting to corrupted supervisory signals.
V-F Effectiveness Analysis of STAND
To rigorously validate the architectural design of the STAND framework, we conduct a two-part ablation study: (1) assessing the necessity of the core components, namely the Temporal Encoding Module (TEM) and the bidirectional hidden state (Bidir.), and (2) evaluating the specific choice of the sequential encoder.
V-F1 Component Ablation
Table V presents the ablation results regarding the TEM and Bidir. across five datasets. Removing the TEM effectively reduces the model to a pointwise MLP, resulting in a significant performance degradation across all metrics (e.g., a $37.2\%$ drop in average score). This decline emphasizes that relying solely on pointwise features is insufficient for time-series anomaly detection; the TEM is indispensable for capturing the sequential context that distinguishes anomalous from normal behavior.
Furthermore, retaining the TEM but disabling Bidir. decreases overall performance by $8.07\%$. This indicates that bidirectional modeling yields critical supplementary information. In time-series analysis, anomalies are often characterized not only by preceding events but also by subsequent recovery patterns. Consequently, Bidir. enables STAND to leverage both historical and future contexts, ensuring more robust timestamp-level classification.
TABLE V: Ablation results of STAND on five datasets. Bidir. with indicates that only the forward hidden state in Eq. (5) is used. TEM with indicates that the Temporal Encoding Module is replaced by an Identity function. Bold and underlining denote the best and second-best performing models, respectively, within a single dataset.
Bidir. TEM Dataset CCE F1 Aff-F1 UAff-F1 AUC V-PR
PSM 9.25 21.08 68.19 42.53 67.37 52.51
38.73 29.93 29.13 28.57 88.44 79.77
46.94 30.33 29.65 29.35 90.74 83.66
SWaT 20.04 43.13 84.81 60.83 76.32 51.76
51.95 58.01 89.12 88.59 90.27 80.63
71.48 57.95 87.84 87.42 92.02 87.14
WADI 14.82 30.07 82.49 53.15 76.65 33.51
60.31 76.48 84.22 74.70 95.07 85.94
65.24 82.03 89.93 84.07 91.72 87.08
Swan 3.14 13.35 32.26 24.52 53.58 88.27
51.08 25.37 33.81 33.54 82.56 97.06
55.53 25.57 46.98 46.65 84.76 97.19
Water 8.92 6.43 88.54 76.25 63.13 6.27
64.04 31.62 83.89 64.44 98.64 84.92
84.44 34.01 96.61 93.41 99.62 95.38
Average 11.23 22.81 71.26 51.46 67.41 46.46
53.22 44.28 64.04 57.97 91.00 85.66
64.73 45.98 70.20 68.18 91.77 90.09
V-F2 Impact of Temporal Encoder Architecture
Having established the necessity of a bidirectional temporal encoder, we further investigate the optimality of the BiLSTM architecture within this supervised setting. We compare the default BiLSTM in STAND against two variants: a Bidirectional GRU (w/ GRU) and a Multi-Head Attention mechanism (w/ Attn).
To quantitatively evaluate the performance differences, we conducted a Wilcoxon signed-rank test across the five datasets. As detailed in Table VI, the BiLSTM-based STAND demonstrates statistically significant improvements over both the Attention and GRU variants in terms of the CCE metric ($p<0.05$). Additionally, STAND significantly outperforms the GRU variant on the VUS-PR metric ($p<0.05$).
These statistical findings validate our architectural choice. The inferior performance of the more complex Attention variant suggests that, under conditions of limited supervision, highly parameterized attention mechanisms may be susceptible to overfitting and offer diminishing returns. Conversely, the GRU variant exhibits insufficient representational capacity to consistently localize complex anomalous events, as reflected by the decreased V-PR scores. Ultimately, the BiLSTM architecture provides an optimal balance, delivering stable structural consistency and superior local anomaly detection capabilities without introducing excessive model complexity.
TABLE VI: Performance comparison of different encoder variants across five datasets. Bold indicates the best performance. ‘*‘ and $\dagger$‘ indicate that STAND is statistically significantly better than the w/ Attn and w/ GRU variants, respectively, at the $p<0.05$ level (using the Wilcoxon signed-rank test).
Variant Dataset CCE F1 Aff-F1 UAff-F1 AUC V-PR
w/ Attn PSM 37.46 30.46 17.03 16.96 83.10 71.54
w/ GRU 35.50 30.34 31.80 31.21 85.10 72.63
STAND 46.94 30.33 29.65 29.35 90.74 83.66
w/ Attn SWaT 69.56 58.13 78.21 77.02 92.08 88.08
w/ GRU 52.31 57.94 81.10 80.92 91.29 79.40
STAND 71.48 57.95 87.84 87.42 92.02 87.14
w/ Attn WADI 61.64 77.55 91.17 82.99 93.39 83.22
w/ GRU 60.48 68.07 89.28 82.56 95.20 72.43
STAND 65.24 82.03 89.93 84.07 91.72 87.08
w/ Attn Swan 54.93 24.97 45.68 45.20 87.37 95.58
w/ GRU 50.22 24.76 42.41 41.90 86.07 95.11
STAND 55.53 25.57 46.98 46.65 84.76 97.19
w/ Attn Water 82.28 28.14 98.14 96.29 97.00 24.33
w/ GRU 83.14 32.15 94.90 89.83 98.51 88.71
STAND 84.44 34.01 96.61 93.41 99.62 95.38
w/ Attn Average 61.17 43.85 66.05 63.69 90.59 72.55
w/ GRU 56.33 42.65 67.90 65.28 91.23 81.66
STAND 64.73∗† 45.98 70.20 68.18 91.77 90.09
$p$-value (STAND vs. w/ Attn) 0.043 0.224 0.345 0.224 0.685 0.079
$p$-value (STAND vs. w/ GRU) 0.043 0.079 0.224 0.138 0.892 0.043
Figure 5: Sensitivity analysis of model performance with respect to three key hyperparameters on five datasets. (a)-(b) show the impact of model dimension ($d_{\text{model}}$) on CCE and VUS-PR, respectively; (c)-(d) illustrate the effect of the number of TEM layers; (e)-(f) depict the influence of window size.
Figure 6: Sensitivity analysis of model performance with respect to three key hyperparameters on three subsets of the PSM dataset. (a)-(b) show the impact of model dimension ($d_{\text{model}}$) on CCE and VUS-PR, respectively; (c)-(d) illustrate the effect of the number of TEM layers; (e)-(f) depict the influence of window size.
V-G Sensitivity Analysis of STAND
To investigate the influence of hyperparameters on STAND, we search for the model’s CCE and VUS-PR performance across various parameter settings on all datasets and the PSM subsets. The results are presented in Figure 5 and Figure 6, respectively, where the error bars represent the 95% confidence interval (all metric results are provided in Appendix C). Overall, the parameter sensitivity of STAND remains largely consistent, regardless of the dataset or the anomaly ratio within the dataset.
Based on Figure 5(a-b) and Figure 6(a-b), increasing the model dimension ($d_{\text{model}}$) effectively enhances the model’s detection consistency (CCE) and local detection capability (VUS-PR) across different datasets.
From Figure 5(c-d) and Figure 6(c-d), it can be observed that increasing the number of model layers is only effective when the data volume is relatively small (e.g., Swan, SP1). For most scenarios, selecting a single layer or a two-layer model yields better results. Similarly, as shown in Figure 5(e-f) and Figure 6(e-f), a window size of $32$ performs distinctly better than a window size of $128$.
Therefore, overall, learning the relationships between different channels in a high-dimensional space is more effective than deepening the model or enlarging the temporal window size.
V-H Visualization Analysis
Figure 7: The performance of competitive models on the entire SWaT test dataset.
Figure 8: Visualization results for the initial segment (time steps $150,000$ to $200,000$) of the partitioned SWaT dataset.
Figure 9: Visualization results for the latter segment (time steps $300,000$ to $350,000$) of the partitioned SWaT dataset.
To further analyze the advantages of STAD methods in terms of practical utility in real-world scenarios, we selected and visualized the outputs of two better-performing UTAD-II models (LSTM and M2N2) along with three STAD models (ExtraTrees, LightGBM, and STAND) on the SWaT dataset.
In Figure 7, the first two rows represent the time series data for two selected channels, while the subsequent rows show the anomaly scores output by the different models. Red highlighting indicates the ground truth anomaly regions, and the red dashed line signifies the division point for the STAD training set.
Figures 8 and 9 are localized visualizations of Figure 7, depicting two key regions corresponding to the left and right sides of the red dashed line, respectively. The meaning of each subplot within these figures is consistent with that of Figure 7.
Overall, the average performance scores for LSTM and M2N2 on SWaT are $68.63$ and $33.29$, respectively. However, in practice, only the long-duration anomaly segment around time step $250,000$ is predicted well. Anomalies in other locations are almost entirely missed. In contrast, all three STAD models successfully predict the majority of anomalies within the training set portion. Crucially, they are also capable of detecting partial anomalies in the unseen test set. Among them, STAND exhibits the lowest rate of false positives, demonstrating superior practical availability. Conversely, the two UTAD-II models fail to predict the anomalies in the latter half of the series.
In the preceding segment (Figure 8), both unsupervised methods (LSTM and M2N2) only output extremely small anomaly scores. This suggests that their anomaly scores are difficult for human interpretation and are prone to causing false negatives (missed detections). Conversely, under the influence of the supervisory signal, the three STAD methods are capable of perfectly predicting the anomalies they have already encountered.
In the subsequent segment (Figure 9), both UTAD methods completely fail to detect the anomalies. Among the STAD approaches, ExtraTrees and STAND successfully detect two anomalous segments, while LightGBM only detects one. However, ExtraTrees and LightGBM also suffer from significant false positives, making it difficult for humans to correctly interpret and locate the true anomalies. In contrast, STAND predicts anomaly segments that are cleaner and more interpretable.
Generally, the UTAD methods did not perform as anticipated; their practical performance on the majority of anomalies is poor. In stark contrast, the outputs of the STAD methods are more reliable and readily interpretable by humans.
VI Conclusion
This paper critically re-evaluates the prevailing research paradigm in time series anomaly detection (TSAD), which heavily favors complex model architectures to address the lack of labels. By proposing STAND, a simple yet effective supervised baseline, we quantify the impact of supervisory signals against architectural sophistication. Our empirical results conclusively show that labels matter more than models. Even with a minimal labeling budget (e.g., 10% of data), STAND significantly outperforms existing complex unsupervised methods in terms of detection accuracy, F1 score, and prediction consistency. These findings reveal the limitations of the current algorithm-centric path and highlight the immense value of limited supervision. Consequently, we suggest that future TSAD research should pivot towards data-centric approaches that maximize the utility of available labels to enhance practical deployment performance. In the future, we plan to extend this data-centric paradigm to distributed big data frameworks and explore active learning strategies to select the most valuable labels from massive data streams efficiently.
Acknowledgments
This work was supported in part by National Natural Science Foundation of China No. 92467109, U21A20478, National Key R&D Program of China 2023YFA1011601, and the Major Key Project of PCL, China under Grant PCL2025A11 and PCL2025A13.
Appendix
This is the appendix of Labels Matter More Than Models: Rethinking the Unsupervised Paradigm in Time Series Anomaly Detection.
A Full Results of Unsupervised Methods
Table A7 presents the complete results for all unsupervised methods (including both UTAD-I and UTAD-II categories).
TABLE A7: Performance comparison of all unsupervised methods across five datasets and six metrics. Bold indicates the best performance, and underlining indicates the second-best performance. Methods highlighted in gray are UTAD-I, and methods highlighted in green are UTAD-II.
Dataset PSM SWaT WADI
Metric CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR
Random -0.83 8.51 67.99 5.53 50.10 32.99 -0.48 7.07 68.96 -3.13 49.96 12.77 -0.07 4.90 69.16 -6.51 49.80 8.47
IForest -5.40 3.75 49.60 -1.69 46.76 29.96 -2.02 6.02 62.11 -9.86 32.74 10.37 7.83 10.99 61.99 -11.06 74.70 17.17
LOF 0.39 10.09 73.93 23.52 51.71 35.40 0.06 8.08 72.90 17.03 50.60 13.68 0.93 9.32 74.42 28.97 55.18 11.16
PCA 6.56 18.77 57.93 52.38 74.13 55.01 4.52 48.84 66.17 30.80 89.73 61.73 14.39 39.07 81.74 58.92 81.07 37.04
HBOS -3.44 8.52 44.79 -18.88 49.90 33.23 9.65 10.38 71.49 24.54 75.08 26.34 9.44 9.86 67.40 6.38 73.97 18.50
KNN -1.78 9.08 70.11 31.22 37.39 29.49 -1.95 4.17 70.06 10.90 14.49 7.78 -0.37 5.23 68.68 -6.21 51.76 10.31
KMeans 0.12 7.85 60.05 44.55 38.64 30.14 4.79 19.49 78.12 46.24 33.06 18.18 6.46 12.72 66.56 28.87 64.70 13.79
OCSVM 2.84 17.03 43.42 30.49 52.63 38.46 10.10 4.77 39.06 21.13 77.36 24.78 13.37 32.12 70.40 47.82 81.78 32.00
AE 7.78 5.19 56.50 51.07 56.50 34.90 16.25 17.87 74.95 40.34 84.69 45.08 15.04 27.00 74.26 42.28 64.38 16.65
CNN 0.57 26.33 69.16 60.48 61.26 52.52 29.04 58.34 5.46 5.46 86.77 60.84 2.39 24.20 78.57 43.04 66.08 24.00
LSTM 0.60 26.86 66.48 57.44 63.88 51.85 29.54 58.32 5.41 5.41 86.63 68.63 2.47 23.98 78.70 43.07 70.80 25.67
TranAD 0.23 20.09 63.50 58.40 60.54 46.85 29.78 58.34 5.34 5.34 88.47 63.50 2.60 23.23 77.80 40.01 72.76 30.27
USAD 1.49 21.39 46.14 44.90 61.53 46.62 29.79 58.34 5.33 5.33 88.71 63.04 8.36 23.01 77.83 39.86 73.60 28.02
Omni 3.93 21.78 23.44 23.04 64.11 44.55 29.80 58.34 5.31 5.31 88.85 59.30 10.84 39.02 72.07 52.59 79.13 38.82
A.T. 1.02 18.20 48.73 42.56 59.13 43.46 29.31 58.34 5.35 5.35 74.39 47.13 3.88 28.02 74.06 22.00 61.97 20.27
TimesNet 0.01 15.34 73.81 35.76 57.19 46.33 0.12 8.77 74.09 30.50 30.16 11.71 0.59 16.76 73.85 18.57 70.07 20.66
M2N2 -0.07 15.26 79.11 47.11 63.35 48.17 0.01 23.99 75.65 42.15 81.15 33.29 1.69 20.75 74.89 22.97 64.38 22.25
LFTSAD -0.32 9.27 63.57 16.99 44.32 31.62 0.81 19.79 65.12 3.58 66.36 24.21 0.28 9.54 53.74 -27.72 48.41 11.49
CATCH 0.62 16.85 84.85 62.92 60.65 51.99 0.02 9.33 73.41 23.84 29.78 11.87 0.26 13.64 74.92 31.67 69.76 21.18
Dataset Swan Water Average
Metric CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR CCE F1 Aff-F1 UAff-F1 AUC-ROC VUS-PR
Random -0.27 8.65 26.68 -0.69 49.79 83.90 0.72 1.53 63.75 -3.53 50.32 3.57 -0.18 6.13 59.31 -1.67 49.99 28.34
IForest 5.85 18.18 11.56 5.15 66.30 85.11 12.35 8.11 60.59 19.21 79.10 10.09 3.72 9.41 49.17 0.35 59.92 30.54
LOF 0.14 10.32 21.67 5.50 49.14 83.10 0.07 2.91 64.47 1.14 50.30 4.33 0.32 8.14 61.48 15.23 51.39 29.53
PCA 1.55 24.75 2.21 2.00 59.62 93.45 11.38 18.03 62.58 47.28 90.68 17.92 7.68 29.89 54.12 38.28 79.05 53.03
HBOS 15.92 18.01 23.08 20.14 78.82 88.77 16.45 8.16 65.15 9.51 82.93 11.87 9.60 10.99 54.38 8.34 72.14 35.74
KNN -3.17 6.67 20.69 -15.08 32.99 83.35 3.40 5.06 65.87 0.35 55.71 8.86 -0.77 6.04 59.08 4.23 38.47 27.96
KMeans -7.41 6.43 11.18 6.27 30.04 81.24 28.38 17.89 72.45 46.33 92.85 24.01 6.47 12.88 57.67 34.45 51.86 33.47
OCSVM 11.82 10.43 10.96 8.70 71.10 85.26 20.57 17.79 70.93 20.01 92.31 22.78 11.74 16.43 46.95 25.63 75.04 40.66
AE 5.95 18.91 21.89 20.80 60.01 88.57 1.76 5.68 54.44 22.81 50.71 2.36 9.36 14.93 56.41 35.46 63.26 37.51
CNN 0.56 24.84 15.92 15.78 72.99 93.71 2.62 3.96 44.78 9.19 54.04 2.55 7.04 27.53 42.78 26.79 68.23 46.72
LSTM 0.38 23.98 27.41 27.11 75.24 92.63 2.62 8.63 54.09 12.65 60.35 5.84 7.12 28.36 46.42 29.14 71.38 48.92
TranAD 0.08 15.94 18.57 5.76 65.01 91.96 0.32 6.20 50.82 13.96 52.11 6.51 6.60 24.76 43.21 24.69 67.78 47.82
USAD 0.54 20.85 11.95 11.42 61.81 92.27 3.21 9.35 41.21 13.84 71.66 6.27 8.68 26.59 36.49 23.07 71.46 47.25
Omni 1.86 24.44 0.26 -0.25 68.55 94.25 3.23 9.30 41.19 5.74 74.37 6.44 9.93 30.58 28.45 17.29 75.00 48.67
A.T. 0.02 3.45 6.64 -5.31 50.27 84.22 1.91 3.29 71.44 30.10 65.94 5.02 7.23 22.26 41.24 18.94 62.34 40.02
TimesNet 0.40 18.79 12.21 3.41 58.88 92.56 0.25 10.11 63.76 -4.59 71.46 19.50 0.28 13.95 59.54 16.73 57.55 38.15
M2N2 0.42 26.50 0.03 0.03 77.54 96.62 0.65 22.13 84.87 63.64 91.91 37.77 0.54 21.73 62.91 35.18 75.67 47.62
LFTSAD 0.17 10.73 15.32 3.05 51.44 84.95 0.47 4.91 69.24 8.06 45.73 6.68 0.28 10.85 53.40 0.79 51.25 31.79
CATCH 0.22 25.60 3.05 3.03 61.55 94.17 0.88 18.98 81.54 54.51 91.28 27.36 0.40 16.88 63.55 35.19 62.61 41.31
Figure A10: Sensitivity analysis of $d_{\text{model}}$ across 5 datasets on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A11: Sensitivity analysis of number of TEM layers across 5 datasets on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A12: Sensitivity analysis of window size across 5 datasets on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A13: Sensitivity analysis of model dimension $d_{\text{model}}$ across 4 subsets of the PSM dataset on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A14: Sensitivity analysis of number of TEM layers across 4 subsets of the PSM dataset on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
Figure A15: Sensitivity analysis of window size across 4 subsets of the PSM dataset on six key performance metrics. (a) CCE; (b) F1 score; (c) Aff-F1 score (Aff-F1); (d) UAff-F1 score; (e) AUC-ROC; (f) VUS-PR.
B Temporal Split Validation Results
Since the supervised methods utilized the initial segment of the original test set as their training data, it is challenging to fairly analyze whether STAD methods genuinely outperform UTAD methods. To mitigate this drawback, we designate the latter segment of the original test set (after partitioning) as the true evaluation set. Both UTAD and STAD methods are tested on this specific set, and the results are presented in Table A8.
Under this fair comparison setting, the best-performing UTAD-I method is PCA, with an average score of $44.34$. For UTAD-II, the best method is LSTM, achieving an average score of $47.09$. In contrast, recent state-of-the-art UTAD-II methods, M2N2, LFTSAD, and CATCH, yield average scores of $37.91$, $27.96$, and $35.71$, respectively. The two best-performing STAD methods are RF and STAND, with average scores of $45.79$ and $54.68$.
We observe that two simple STAD methods outperform the majority of UTAD methods, and the performance of STAND is significantly superior to both other method categories. Furthermore, we note that recent UTAD-II methods (M2N2, LFTSAD, and CATCH) are surprisingly less competitive than older, simpler approaches (PCA and LSTM). In addition, the prediction consistency (CCE) performance of all UTAD methods is poor, falling far behind that of the STAD approaches. This suggests their limited practical applicability.
TABLE A8: Three method categories’ performance comparison on the latter half of the divided dataset (across five datasets and six metrics). Bold indicates the best performance, and underlining indicates the second-best performance. Methods highlighted in gray are UTAD-I, methods highlighted in green are UTAD-II, and methods highlighted in orange represent STAD methods.
Dataset PSM SWaT WADI
Metric CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR
Random -0.52 8.49 71.43 11.66 50.20 35.30 -0.46 7.87 69.35 -4.59 50.06 19.55 0.99 3.18 66.37 20.87 51.02 6.17
IForest 0.40 14.23 49.22 -1.25 51.34 36.30 0.40 6.72 67.58 -3.93 32.17 15.36 11.57 10.59 72.41 26.20 85.22 19.86
LOF 0.84 11.90 77.47 29.69 52.37 39.09 0.07 7.34 72.47 11.62 49.90 19.73 1.16 7.94 73.75 27.64 54.82 8.02
PCA 10.22 17.77 56.73 51.31 78.32 59.91 5.64 34.75 61.50 -2.99 92.11 77.11 16.32 32.83 83.96 58.69 84.93 29.05
HBOS -0.50 10.66 54.69 8.80 48.40 34.71 14.58 5.07 71.07 26.69 85.23 51.92 16.01 18.53 75.88 35.34 85.02 26.56
KNN -1.87 6.01 70.79 38.02 30.14 26.60 -2.94 1.93 67.76 -2.78 9.41 11.14 -0.12 3.04 68.35 -28.87 57.96 8.82
KMeans -1.10 6.79 63.80 53.36 35.30 28.84 0.53 14.31 77.65 40.88 21.37 17.78 13.50 5.43 60.69 0.36 73.85 10.92
OCSVM -0.60 7.66 52.24 41.95 41.82 31.52 10.49 15.86 40.40 19.28 86.41 45.01 20.36 26.60 70.09 38.04 90.25 30.55
AE 5.94 4.90 35.00 14.38 59.18 37.57 21.36 31.37 76.01 46.24 91.72 76.31 22.08 24.35 73.78 45.87 77.90 18.94
CNN 2.06 24.97 49.02 28.51 78.98 62.25 37.30 41.57 11.57 11.57 91.75 82.35 13.53 28.86 86.69 64.58 87.09 35.37
LSTM 1.87 25.54 53.94 37.36 77.35 61.26 38.07 41.57 12.29 12.29 89.88 81.90 14.65 29.25 86.66 66.06 83.47 35.49
TranAD 0.47 16.40 47.32 43.45 69.66 53.91 38.32 41.57 12.26 12.26 92.69 81.38 12.60 27.93 85.20 66.75 84.60 36.17
USAD 4.28 16.72 44.31 40.91 72.39 55.40 38.30 41.57 12.23 12.23 92.40 80.85 15.29 30.31 86.20 67.15 76.46 29.52
Omni 11.96 18.16 17.55 16.05 72.61 51.57 38.41 41.57 11.16 11.16 92.78 77.46 14.66 30.18 79.33 46.35 78.19 42.14
A.T. 0.51 10.26 63.57 -3.85 50.65 36.48 0.00 1.20 65.24 -14.25 49.26 19.53 -0.34 1.06 35.84 -53.36 36.60 5.76
TimesNet -0.01 12.76 74.16 27.57 55.98 45.34 0.05 4.39 71.47 7.50 19.79 13.50 2.18 15.49 72.32 24.23 70.38 21.66
M2N2 -0.01 14.83 80.84 46.18 65.93 51.71 -0.51 15.32 27.56 25.80 80.08 40.06 10.63 27.13 81.07 38.19 83.14 35.44
LFTSAD 1.02 18.75 63.38 28.61 56.44 44.73 0.73 22.53 70.22 3.36 73.46 45.84 3.30 14.82 69.27 34.03 57.60 13.61
CATCH 0.54 15.26 83.83 56.61 61.46 50.61 0.01 5.79 72.34 10.56 19.74 13.57 2.23 15.75 76.92 39.90 80.45 26.05
RF 13.14 18.81 50.78 21.89 73.17 50.54 31.01 39.04 60.07 56.87 90.84 80.28 23.82 46.73 73.04 60.38 85.49 63.63
SVM 12.25 22.13 65.87 52.31 80.42 57.39 19.93 15.09 68.89 25.96 91.67 76.17 12.87 33.22 81.05 50.59 81.50 42.92
AdaBoost 10.05 20.06 34.14 28.86 70.80 39.94 15.39 22.34 47.91 34.16 84.07 37.79 11.60 27.13 58.41 24.08 78.62 27.07
ExtraTrees 10.78 20.36 53.19 36.77 79.11 57.64 39.10 52.71 12.91 12.89 88.88 53.93 27.43 51.47 88.03 70.48 94.61 73.56
LightGBM 6.79 16.35 55.07 16.70 71.52 51.89 39.25 28.39 30.33 -12.15 89.38 64.47 23.50 44.47 77.52 50.19 86.81 50.02
STAND 16.39 23.58 64.18 55.61 66.81 55.47 51.97 36.88 76.94 69.73 90.68 79.62 26.70 37.33 78.10 55.51 73.17 39.71
Dataset Swan Water Average
Metric CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR CCE F1 Aff-F1 UAff-F1 AUC V-PR
Random -0.10 9.10 26.79 4.24 49.71 89.09 4.74 0.55 66.75 0.38 55.06 0.49 0.93 5.84 60.14 6.51 51.21 30.12
IForest 5.77 17.29 3.23 -3.03 68.63 92.47 37.09 8.07 52.82 -47.52 98.97 17.21 11.05 11.38 49.05 -5.91 67.27 36.24
LOF 0.05 10.84 23.54 14.35 48.62 90.63 6.58 2.10 67.15 0.60 85.45 1.86 1.74 8.02 62.88 16.78 58.23 31.87
PCA 1.52 18.20 1.35 1.33 53.33 96.02 46.62 7.74 77.31 41.62 99.01 37.89 16.06 22.26 56.17 29.99 81.54 60.00
HBOS 13.70 15.04 18.22 13.47 80.02 96.18 31.77 6.85 56.58 -39.18 96.26 9.63 15.11 11.23 55.29 9.02 78.99 43.80
KNN -3.79 7.71 22.44 -15.13 31.37 88.35 22.65 4.20 49.15 -52.68 70.61 5.60 2.79 4.58 55.70 -12.29 39.90 28.10
KMeans -9.32 4.74 16.58 8.88 23.50 87.55 44.24 6.19 39.83 -66.82 93.74 24.05 9.57 7.49 51.71 7.33 49.55 33.83
OCSVM 11.64 11.31 12.23 10.84 68.59 92.06 37.64 8.40 75.02 34.46 99.35 21.63 15.91 13.97 50.00 28.91 77.28 44.15
AE 13.97 13.59 21.21 20.46 68.29 92.43 17.22 1.88 71.53 22.88 58.07 2.88 16.11 15.22 55.51 29.97 71.03 45.63
CNN 0.69 18.42 0.08 0.08 66.83 96.12 31.41 8.51 90.19 78.80 99.87 59.99 17.00 24.47 47.51 36.71 84.90 67.22
LSTM 0.50 16.37 29.49 29.11 71.67 95.29 27.28 5.31 93.18 85.62 91.41 18.73 16.47 23.61 55.11 46.09 82.76 58.53
TranAD 0.08 11.20 19.84 -7.58 60.35 94.03 19.12 5.20 91.42 81.54 94.48 20.85 14.12 20.46 51.21 39.28 80.35 57.27
USAD 0.70 16.44 3.97 3.81 56.80 95.05 29.06 5.53 90.52 79.71 93.95 20.92 17.53 22.11 47.45 40.76 78.40 56.35
Omni 2.45 18.34 0.08 0.08 59.54 95.63 18.24 3.98 92.22 83.31 92.90 42.93 17.14 22.45 40.07 31.39 79.20 61.95
A.T. 0.01 5.64 18.16 -12.74 49.19 89.65 15.94 2.99 51.39 -43.74 61.45 2.99 3.22 4.23 46.84 -25.59 49.43 30.88
TimesNet 0.59 18.48 0.08 0.08 53.38 94.57 0.98 3.76 67.92 10.57 82.65 25.59 0.76 10.97 57.19 13.99 56.44 40.13
M2N2 0.53 18.45 0.08 0.08 85.05 98.41 7.69 5.64 70.36 13.77 91.45 22.53 3.67 16.27 51.98 24.80 81.13 49.63
LFTSAD 0.09 12.80 15.11 -5.85 44.97 90.26 0.11 1.88 51.36 -46.26 51.50 1.12 1.05 14.16 53.87 2.77 56.79 39.11
CATCH 0.40 18.48 0.08 0.08 55.82 95.67 22.06 5.53 75.30 32.23 89.20 44.83 5.05 12.16 61.69 27.88 61.33 46.15
RF 27.51 16.68 29.91 28.55 87.60 95.32 53.80 19.66 65.81 60.20 87.16 26.33 26.14 17.93 54.32 40.44 84.07 51.85
SVM 12.31 16.55 25.65 24.73 82.46 97.19 19.17 7.24 66.59 42.27 65.06 12.07 11.31 12.00 51.72 33.02 77.30 47.20
AdaBoost 33.24 20.19 27.23 25.98 87.31 93.06 33.52 17.43 98.42 96.84 88.12 32.65 17.59 14.29 46.72 28.74 78.36 45.24
ExtraTrees 27.90 17.86 28.26 26.99 87.35 95.54 26.57 13.18 97.91 96.04 85.31 11.49 25.52 18.39 50.63 35.95 84.59 52.82
LightGBM 38.09 16.29 36.29 35.21 89.54 94.90 21.21 26.39 33.03 -2.69 80.66 3.80 24.28 14.79 49.23 21.57 80.89 47.88
STAND 33.26 17.17 27.39 26.95 77.27 96.69 64.19 6.63 92.21 83.33 96.53 20.37 38.50 24.32 67.76 58.22 80.89 58.37
C Full Results of STAND Sensitivity Analysis
To investigate the parameter sensitivity of STAND, we conducted a comprehensive parameter analysis across all five real-world datasets and the four PSM subsets.
Figures A10, A11, and A12 illustrate the impact of the model dimension ($d_{\text{model}}$), the number of TEM layers, and the window size on STAND’s performance across the five datasets, respectively. Overall, the model’s performance improves (for almost all metrics) as the model dimension increases. Conversely, performance tends to degrade (for both point-wise and event-level metrics) as the number of TEM layers and the temporal window size increase.
Figures A10, A11, and A12 also present the impact of $d_{\text{model}}$, the number of TEM layers, and the window size on STAND’s performance across the four PSM subsets. Generally, the trend of model performance change is consistent with that observed on the five real-world datasets. Specifically, when the model dimension increases, the performance improves on three subsets (SP2–SP4), while the performance change on SP1 is relatively marginal. Conversely, when the number of TEM layers and the temporal window size increase, performance slightly decreases on the SP2–SP4 subsets, but no clear trend of performance change is observed on SP1.
Therefore, the influence of hyperparameters on STAND is primarily manifested as follows: When the dataset size is large, the impact is significant, exhibiting a clear performance trend; when the data volume is small, the model is more severely affected by random initialization, leading to less discernible performance trends.
Furthermore, it is important to note that the average level of Aff-F1 and UAff-F1 scores is noticeably better when the training set is SP1 compared to SP2–SP4. This is because these two metrics require the setting of an anomaly evaluation buffer, where predictions within the buffer zone are also regarded as correct. When the anomaly ratio is low, this type of tolerance metric can easily lead to inaccurate evaluations.